### v1

# Trading Carbon for Physics: On the Resource Efficiency of Machine Learning for Spatio-Temporal Forecasting

Abstract
Development of modern deep learning methods has been driven primarily by the push for improving model efficacy (accuracy metrics). This sole focus on efficacy has steered development of large-scale models that require massive resources, and results in considerable carbon footprint across the model life-cycle. In this work, we explore how physics inductive biases can offer useful trade-offs between model efficacy and model efficiency (compute, energy, and carbon). We study a variety of models for spatio-temporal forecasting, a task governed by physical laws and well-suited for exploring different levels of physics inductive bias. We show that embedding physics inductive biases into the model design can yield substantial efficiency gains while retaining or even improving efficacy for the tasks under consideration. In addition to using standard physics-informed spatio-temporal models, we demonstrate the usefulness of more recent models like flow matching as a general purpose method for spatio-temporal forecasting. Our experiments show that incorporating physics inductive biases offer a principled way to improve the efficiency and reduce the carbon footprint of machine learning models. We argue that model efficiency, along with model efficacy, should become a core consideration driving machine learning model development and deployment. [^1] Source code for the experiments in this work are available at: https://github.com/sophiawilson18/FlowMatching.
1 Introduction
Machine learning (ML) has advanced rapidly over the past decade (Sevilla et al., 2022), driven largely by the pursuit of improving model efficacy (accuracy metrics). This narrow focus has led to increasingly large models that require substantial computational resources and generate significant carbon emissions across their lifecycle (Strubell et al., 2019; Anthony et al., 2020; Luccioni et al., 2023). As models scale, efficacy gains diminish while costs escalate, underscoring the need to move beyond accuracy as the sole measure. A more holistic assessment must balance efficacy with efficiency, capturing not only predictive performance but also compute, energy, and carbon costs (Henderson et al., 2022). This requires exploring trade-offs and identifying strategies that strike a favourable balance, rather than lopsided optimization.
Figure 1: a: Conceptual map of modelling approaches by data availability (vertical axis) and domain knowledge (horizontal axis). Physics-informed ML lies in the middle, combining data-driven flexibility with physics priors. b: Test MSE versus CO2eq emissions for a data-driven NN (red squares), an unsupervised PINN (blue dots), and a semi-supervised PINN (grey triangles). The dashed line marks the Pareto front, dominated by the S-PINN, indicating the best accuracy-carbon trade-off. Details of the experiment are in Sec. 4.2.
Physics-informed ML offers a promising way to achieve this balance. Physics priors such as conservations laws, symmetries, or governing equations can be embedded into model design through, e.g., loss terms, architectural layers, or mechanisms. This combines the flexibility of data-driven methods with the structure of domain knowledge (Fig.1-a). As shown in Fig. 1-b, a semi-supervised physics-informed neural network (PINN) achieves the best accuracy–carbon trade-off, outperforming both a purely data-driven neural network (NN) and a physics-only PINN (see Sec. 4.1 for details). This example highlights how hybrid approaches can yield superior efficiency without sacrificing efficacy.
Building on this intuition, we extend the perspective from task-specific models such as PINNs to general-purpose architectures that embed physics inductive biases in ways that are transferable across datasets and domains. We focus on spatio-temporal forecasting of partial differential equations (PDE) dynamics, a setting characterised by high computational cost and practical relevance. Operational weather and climate forecasting, for instance, requires large-scale PDE solvers running on supercomputers multiple times per day (Lam et al., 2023). In such contexts, even modest efficiency improvements can translate into substantial operational carbon savings, making this domain an ideal test-bed for exploring sustainable ML design (Van Wynsberghe, 2021). Our key contributions towards this end are:
  1. 1.
    Carbon-aware evaluation: We promote evaluation practices that consider both predictive accuracy and carbon cost, explicitly characterising trade-offs.
  2. 2.
    Characterising the accuracy–carbon trade-off: We conduct three experiments on PDE dynamics and spatio-temporal forecasting to compare a range of models with varying levels of physics inductive bias. Our main study focuses on forecasting incompressible shear flow, comparing U-nets (no/weak bias), flow matching (medium bias), and Fourier neural operators (strong bias).
Figure 2: Modelling the harmonic oscillator. Training and extrapolation predictions when using NNs with different activation functions: ReLU, Tanh, Sine, Snake, and using a PINN.
2 Background and Related Work
Revisiting the Bias-Variance Trade-off.
It is commonly understood in ML that strong model assumptions (bias) risk under-fitting, whereas high sensitivity to training data can lead to over-fitting (variance) (Kohavi et al., 1996). Striking the right balance is crucial to obtain the right class of models.
The same trade-off can also offer a useful way to perform model selection when viewed through the lens of resource consumption. Typically, if domain knowledge (physics-based or others) can be incorporated into models as inductive bias, it can reduce the reliance on training data or reduce prediction errors. This is illustrated in Fig. 2 for the task of modelling a harmonic oscillator. Models with appropriate inductive bias (sinusoidal oscillations) perform better for the same number of training data points and model parameters. See Sec. 4.1 for more details.
The Question of Resource-Awareness.
The standard metrics used to characterise the performance of ML models primarily focus on efficacy and are untethered from efficiency considerations. This introduces a preference for models that could offer high efficacy at large resource costs (Bakhtiarifard et al., 2024). There are two primary streams of work that aim to include resource-aware metrics: using composite metrics and multi-objective optimisation.
Several composite metrics have been proposed to combine predictive performance with resources like energy consumption or carbon footprint. Evchenko et al. (2021) introduced a composite metric that combines efficacy and efficiency metrics to study the influence of resource constraints on traditional (non deep learning) ML models. Carburacy (Moro et al., 2023) extends the idea to transformer models by jointly quantifying accuracy and carbon emissions. More recently, Kapoor et al. (2025) proposed EcoL2, a composite metric similar to Carburacy but extended to include the costs of data generation and hyperparameter tuning.
Multi-objective optimisation has been primarily used in neural architecture search (NAS) resulting in models like EfficientNet (Tan and Le, 2019). More recently, energy consumption (Bakhtiarifard et al., 2024) and carbon footprint (Zhao et al., 2024) have been taken into consideration in NAS. These methods, however, have not actively studied the influence of physics inductive biases on the resource consumption of ML models.
Spatio-Temporal Forecasting Models.
There are many model classes for the task of spatio-temporal forecasting; each differing in the extent of physics inductive bias they use. Fig. 3 provides a high-level overview of this span for models considered in this work.
Figure 3: Spectrum of model families from purely data-driven ML to physics-based solvers.
U-net. U-nets (Ronneberger et al., 2015) can be used for spatio-temporal forecasting by mapping a window of past states to future ones. They capture local correlations by fusing features across multiple spatial scales, but do not explicitly constrain the dynamics beyond what is enforced by the training data. Mild forms of inductive biases can be introduced to convolutional architectures like U-net by including circular padding to respect periodic boundaries and avoid edge discontinuities.
Flow Matching. Recent classes of probabilistic generative models like flow matching (FM) (Lipman et al., 2023) can be viewed as an intermediate class of models that can include some degree of inductive bias by modelling dynamics as continuous-time flows. This formulation resonates with dynamical systems and intuitively suggests smoother trajectories. FM has been applied to spatio-temporal forecasting and other PDE dynamics tasks (Lim et al., 2025; Li et al., 2025), with Baldan et al. (2025) further embedding PDE residuals into the objective.
Fourier Neural Operator. The Fourier Neural Operator (FNO) (Kovachki et al., 2023) introduces a physics prior through its Fourier-space representation, enabling the model to capture global spectral interactions. This makes it particularly well-suited to systems with periodicity and long-range correlations. In our spatio-temporal forecasting setting, we adapt the FNO from its original operator-learning formulation (Lu et al., 2021) to an autoregressive setup, where its spectral representation provides a strong inductive bias compared to the U-net by explicitly encoding global and periodic structure.
Physics-Informed Neural Network. PINNs (Raissi et al., 2019) embed governing PDEs directly into the training objective, thereby regularising the optimisation landscape and guiding solutions towards physical consistency. By minimizing PDE residuals, PINNs can even be trained without supervised data, representing the strongest inductive bias among the models considered. Embedding conservation laws, symmetries, or governing PDEs into the ML model development pipeline has been shown to reduce the number of trainable parameters (Dutta et al., ; Patra et al., 2024), reduce the amount of training data (Psaros et al., 2023; Zhong et al., 2021), and accelerate convergence (Brehmer et al., 2024; Jahani-nasab and Bijarchi, 2024).
3 Methods
Problem Setup.
We consider spatio-temporal forecasting of dynamical systems in the general setting, where the objective is to approximate the solution $u(x,y,t)$ of a PDE given basic inputs such as initial and boundary conditions. We frame this as an autoregressive prediction task on a two-dimensional, uniformly spaced mesh $(x,y)\in\{x_{1},x_{2},\ldots,x_{H}\}\times\{y_{1},y_{2},\ldots,y_{W}\}$ with discretized time steps $t\in\{t_{1},t_{2},\ldots,t_{T}\}$. At each step $t_{k}$ the state is represented as a tensor $u_{k}(x,y)\in\mathbb{R}^{H\times W\times C}$, where $k\in\{1,2,..,T\}$ indexes discrete time steps in a trajectory of length $T$. $H$ and $W$ are spatial dimensions and $C$ is the number of physical fields (channels), such as pressure or velocity components.
To capture temporal dependencies, the model is conditioned on a finite history of length $h$. The input window is denoted as $u_{k-h+1:k}(x,y):=(u_{k-h+1}(x,y),\ldots,u_{k}(x.y))$. A parameterised function $f_{\theta}$ is trained to predict:
$u_{k+1}(x,y)=f_{\theta}(u_{k-h+1:k}(x,y)\big).$ (1)
During training, the loss is computed on one-step predictions. At inference, we assess long-term stability via autoregressive rollouts, obtained by iteratively applying Eq. 1 and feeding the model’s own predictions back as inputs.
We adapt this generalised model formulation to three problems. Each problem is then tackled with relevant set of methods that range from being entirely data-driven to offering relevant degrees of physics inductive biases.
  1. 1.
    Harmonic Oscillator. Modelling the dynamics of the harmonic oscillator only requires the time variable i.e, $u(t)$. For the harmonic oscillator, Eq. 1 reduces to $u_{k+1}=f_{\theta}(u_{k-h+1:k}\big)$, where the spatial variables $(x,y)$ are dropped.
    We assess the impact of appropriate physics-inductive biases by comparing simple feed-forward NNs with and without biases. Specifically, these biases are introduced through activation functions and a loss term, for modelling the damped and undamped oscillators.
  2. 2.
    Viscous Burgers’ Equation. Solving the viscous Burgers’ equation increases the modelling complexity compared to the harmonic oscillator, as it has one spatial and one temporal variable to be considered. Eq. 1 for this case is given as: $u_{k+1}(x)=f_{\theta}(u_{k-h+1:k}(x)\big)$, where only one spatial variable $x$ is modelled along time.
    We evaluate how different combinations of data and physics constraints influence predictive accuracy and efficiency by comparing a supervised NN, a physics-only PINN, and a hybrid PINN for approximating the viscous Burgers’ equation.
  3. 3.
    Incompressible Shear Flow. Modelling the spatio-temporal dynamics of incompressible shear flow governed by Navier-Stokes equations is the most complex task considered. It has dependence on both the spatial and temporal variables. As a result, we use Eq. 1 in its original formulation.
    We evaluate several models spanning a broad spectrum of physics biases, depicted in Fig. 3. Purely data-driven models like U-net, the moderately biased model FM, and strongly biased neural operators are adapted for this task. The U-nets and FNOs use a history of $h=4$ states, i.e. they are trained for one-step prediction according to Eq. 1 i.e., $u_{k+1}(x,y)=f_{\theta}(u_{k-3:k}(x,y)\big).$
    FM replaces direct one-step prediction with a continuous transport between consecutive states. Instead of learning a map $u_{k-h+1:k}(x,y)\mapsto u_{k+1}(x,y)$, FM predicts the next state by transporting the current state forward along a learned flow. Starting from $u_{k}(x,y)$, the method defines a flow trajectory $\phi_{\tau}(u_{k})$ over a synthetic time axis $\tau\in[0,1]$, where the flow evolves under a time-dependent vector field $v_{\tau}$. The one-step forecast is then obtained as the endpoint of this flow: $\hat{u}_{k+1}(x,y)=\phi_{1}(u_{k}(x,y))$, where $\partial_{\tau}\phi_{\tau}(u)=v_{\tau}(\phi_{\tau}(u)),\phi_{0}(u)=u_{k}(x,y).$ In practice, a NN is trained to approximate this vector field, defined by an interpolation path between consecutive states. Further FM details including the interpolation path, vector field, and learning objective are provided in App. A.
4 Data and Experiments
Based on the problem formulation in Sec. 3, we set-up experiments on the three tasks to explore the trade-off between efficacy and efficiency. Specific choices for each of the experiments are elaborated in detail next.
4.1 Harmonic Oscillator
We use simple feed-forward NNs to model the harmonic oscillator by introducing physics biases in two ways: through activation functions, where Sine and Snake (Ziyin et al., 2020) have periodic biases, and through a physics-informed loss objective. ReLU and Tanh serve as unbiased baselines.
The training data is generated from the damped harmonic oscillator, $u(t)=e^{-\delta t}(2A\cos(\phi+\omega t))$, where: $A=1/(2\cos\phi),\penalty 10000\ \penalty 10000\ \phi=\arctan\left(-\delta/\omega\right)\penalty 10000\ \penalty 10000\ \mathrm{and}\penalty 10000\ \penalty 10000\ \omega=\sqrt{\omega_{0}^{2}-\delta^{2}}\penalty 10000\$, with added Gaussian noise $(\mu=0,\sigma=0.3)$. We consider both the non-damped ($\delta=0$) and the damped ($\delta=0.02$) regimes.
All models use the same underlying architecture (two hidden layers of width 64), trained for $100$ epochs with Adam optimizer (Kingma and Ba, 2015) using a batch size of 25. Learning rates are $0.01$ for standard networks and $0.001$ for the PINN, which employs Snake activations. Results are averaged over $10$ runs. Additional details are provided in App. B.
4.2 Viscous Burgers’ Equation
Figure 4: Target and training data for viscous Burgers’ equation. a: Target data. b: Supervised samples. c: PDE residuals samples. d: Initial and boundary condition points.
In this experiment, we study models to approximate the solution of the viscous Burgers’ equation with Dirichlet boundary conditions on a 2D spatio-temporal grid:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}=0,$
$u(x,0)=-\sin(\pi x),$
$u(1,t)=u(-1,t)=0.$
for $x\in[-1,1]$, $t\in[0,1]$, and with fixed viscosity ($\nu=0.025$). The domain is discretized with $256$ spatial points and $100$ time steps. We compare three models with varying physics inductive biases, all using the same underlying architecture (comprising four hidden layers, width $50$, tanh activation functions):
  • NN: a supervised NN trained with the mean-squared error (MSE) on ground-truth data, using the loss objective: $\mathcal{L}_{\text{NN}}=\mathcal{L}_{\text{data}}$.
  • U-PINN: an unsupervised PINN trained only on physics constraints that minimizes PDE residuals, initial conditions (IC) and boundary conditions (BC). Its loss function, $\mathcal{L}_{\text{U-PINN}}$, is composed of the residuals of the PDE and the deviations from the initial/boundary conditions, expressed as: $\mathcal{L}_{\text{U-PINN}}=\mathcal{L}_{\text{PDE}}+\mathcal{L}_{\text{IC}}+\mathcal{L}_{\text{BC}}$.
  • S-PINN: a semi-supervised PINN where half of the training points are used for PDE residuals and boundary/initial conditions, while the other half enforce a data-driven penalty (MSE). Its loss objective is given as, $\mathcal{L}_{\text{S-PINN}}=\mathcal{L}_{\text{U-PINN}}+\mathcal{L}_{\text{NN}}$.
The specific formulations of all three loss objectives are provided in App. C. All models were trained using Adam optimizer followed by L-BFGS optimizer (Liu and Nocedal, 1989) with learning rate $10^{-3}$. Dataset sizes range from $10^{3}$ to $5\times 10^{3}$ training points for PINNs and up to $8\times 10^{3}$ for the NN. Data and code are adapted from Khoa (2022).
4.3 Incompressible Shear Flow
We set-up the large-scale experiment on modelling incompressible shear flow governed by the Navier–Stokes equations, using the spatio-temporal forecasting models introduced in Section 2.
Data.
The incompreesible shear flow data is derived from The Well dataset (Ohana et al., 2024) generated with Dedalus (Burns et al., 2020). We use a subset of the full dataset consisting of 240 simulations spanning six PDE parameter settings with $40$ initial conditions each. Simulations are resolved at $256\times 512$ with $200$ time steps ($\Delta t=0.1$). The data is split by initial condition into 80/10/10 for training, validation, and testing, yielding $37,632$ one-step samples for training, $4,704$ for one-step inference, and $24$ trajectories for autoregressive rollout (initialised at $t=0$). The dataset is particularly challenging due to the coupled evolution of four physical fields (velocity components, pressure, tracer) and its non-linear spatiotemporal dynamics at high resolution. Further details on PDE formulation and initialization are provided in App. D.
Model(s) Physics inductive bias Strength
UNet, UNet-CN None None
UNet-CP, UNet-CN-CP Circular padding Weak
FM Continuous-time vector flow Medium
FNO, TFNO, UFNO Spectral layers Strong
Table 1: Overview of the eight models and the relative strength of their physics inductive biases.
Models.
We evaluate eight architectures spanning a spectrum of physics inductive biases, from purely data-driven to strongly physics-informed, summarised in Table 1. We use standard U-nets (Ronneberger et al., 2015) and a ConvNeXt variant (UNet-CN) (Liu et al., 2022) for improved hardware-efficiency. Zero-padding serves as our no-bias baseline, while circular padding (CP) introduces a weak prior consistent with periodic boundary conditions in the dataset. We also include our proposed FM model. Finally, we include three variants of the Fourier Neural Operator (FNO): a baseline model, a Tucker-factorized (TFNO) variant (Kossaifi et al., 2023) that improves parameter effiency, and a U-shaped (UFNO) variant (Rahman et al., 2023) designed to better capture multi-scale interactions.
Training Procedure.
Across all models, we use a batch size of 16 and the AdamW optimizer (Loshchilov and Hutter, 2019) with $10^{-4}$ weight decay. For U-nets and FNOs, we adopt the coarse-tuned learning rates reported in The Well ($5\cdot 10^{-4}$ and $10^{-3}$, respectively). For FM, we performed a separate coarse tuning and selected $5\cdot 10^{-3}$. A linear warm-up cosine scheduler was applied for the first three epochs. The loss optimised is MSE averaged across fields and space. Training ran on a single Nvidia A40 GPU ($\leq$ 24h) with early stopping (patience 6). FM was trained in low-resolution data space with subsequent upsampling, avoiding the additional overhead of an autoencoder. Further details on FM in low-resolution data space are provided in App. E, and complete model configurations are listed in App. F.
4.4 Energy and Carbon Metrics
We employ Carbontracker (Anthony et al., 2020) to estimate the energy consumption and carbon footprint during training and inference by monitoring real-time hardware power usage (CPU, GPU, DRAM) and adjusting for infrastructure overhead via the power usage effectiveness. The carbon footprint is reported as carbon dioxide equivalent (CO2eq) emissions, which converts all greenhouse gases into an equivalent CO2 warming potential. It is derived from the total energy consumed and by using the global average carbon intensity for 2024 at $445$ gCO2/kWh, ensuring fair comparisons across models (IEA, 2025).
Figure 5: Damped harmonic oscillator predictions. Mean predictions (solid lines) with standard deviation (shaded areas) from $10$ runs. Results are shown for the two best-performing models: Snake and PINN.
5 Results
5.1 Harmonic Oscillator Toy Experiment
The results in Fig. 2 show that models with non-periodic activations perform poorly outside the training domain (extrapolation). Periodic activations (Sine, Snake) offer modest improvements, with the model using Snake activation function performing best. The PINN achieved superior results in both interpolation and extrapolation tasks; most notably for the damped harmonic oscillator shown in Fig 5. It remained stable and accurate even as prediction range increased and dynamics grew more complex.
5.2 Viscous Burgers’ Equation Experiment
For the viscous Burgers’ equation experiment, the comparison with NN, U-PINN, and S-PINN show interesting trends. Both PINNs achieve consistently lower test losses at lower carbon footprints compared to the supervised NN (see Fig. 1-b). Although the NN has the lowest carbon footprint for fixed epochs and data size, achieving a given accuracy is up to $10$ times more carbon-efficient with the PINNs. This S-PINN, which mixes data and physics, dominates the Pareto front in Fig. 1-b, suggesting that combining physics and data gives the model which balances efficacy and efficiency.
5.3 Incompressible Shear Flow Experiment
The first two simple experiments discussed above establish key insights into how physics can affect the model performance and resource consumption for specific tasks. In the incompressible shear flow experiment, we extend this analysis to a more complex task using a wide spectrum of models.
Predictive Performance.
We first assess predictive performance using Pearson correlation $r$ at each step of a 20-step rollout, see Fig. 6. All eight models (see Table 1) achieve high accuracy for the first few steps (r $\approx$ 1), after which performance declines at different rates. Over longer horizons, UNet-CN, UNet-CN-CP, and UFNO emerge as the strongest performers.
Figure 6: Pearson r over 20 rollout steps. Values are averaged over the four fields: velocity components, tracer, and pressure. Models labels are ordered from no bias (top) to strong bias (bottom).
Figure 7 illustrates these stability differences visually. UNet and UNet-CP deteriorate rapidly, accumulating distortions within just a few steps. The modernised variants (UNet-CN, UNet-CN-CP) show marked improvements: UNet-CN captures fine-scale detail but gradually disrupts large-scale dynamics, while UNet-CN-CP better preserves both large and fine-scale structure at the cost of high-frequency artifacts. FM exhibits good temporal stability, though it tends to exaggerate some fine-scale structures. FNO and TFNO maintain the dominant large-scale dynamics but exhibit a systematic drift toward higher values, fail to capture fine-scale variability and eventually introduce distortions. UFNO better recovers both large- and small-scale structures, though these gradually smooth out over time.
Figure 7: Example rollout predictions for one of the four fields (tracer). Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout, illustrating differences in long-horizon stability and fidelity. Models labels are ordered from no bias (top) to strong bias (bottom). Results for the three other fields are provided in Appendix G.
Training Cost.
[…]
Inference Cost.
To provide a more complete evaluation, we also account for inference costs of all the trained models. This can be useful in choosing spatio-temporal forecasting models as inference costs can dominate over training or development costs.
Figure 8: Predictive performance and carbon footprint. a: Predictive accuracy (Pearson r) versus training emissions (kgCO2eq); each point represents every second step of a 20-step rollout (lighter shades indicate earlier steps), averaged over four fields. b: Training emissions (kgCO2eq) from a single run per model. Models are ordered from no bias (left) to strong bias (right). c: One-step inference emissions (gCO2eq) from three repeated runs over the full test set; error bars indicate variability across repetitions. Note that inference costs also include the initial model loading cost.
[…]
Padding choices further illustrate the sensitivity of emissions to architectural design. Switching from zero to circular padding increases training emissions for both UNet and UNet-CN. At inference, the effects diverge: circular padding slightly increases the footprint for UNet, while significantly reducing it for UNet-CN. This highlights how even minor inductive design choices – in this case enforcing circular padding to reflect periodic boundaries – can shift the efficacy–efficiency balance.
Table 6 in App. G reports both training and inference footprints alongside run times. While runtime and energy consumption/ carbon emissions are closely linked, they are not strictly proportional, as some operations could take longer but use less energy, and vice-versa (Henderson et al., 2022; Bakhtiarifard et al., 2024). Additional results for these models can be found in App. G.
6 Discussions
Our study demonstrates that physics inductive biases impact the trade-off between efficacy and efficiency. In simple settings, such as the harmonic oscillator, adding periodic structure or a physics-informed loss enables extrapolation beyond the training range. With viscous Burgers’ equation, combining data with physics yields the model that dominates the Pareto front, achieving lower test errors at lower carbon cost compared to the purely data-driven baseline.
In the incompressible shear flow setting, we observe clear differences in how the physics priors affect both predictive performance and carbon footprint across training and inference. Models with strong physics inductive biases achieve the lowest training footprints, but this comes at the expense of higher inference costs compared to U-net variants. UFNO achieves one of the most favourable overall profiles, combining strong predictive accuracy with a low training footprint. Its inference cost, however, remains nearly twice that of the best-performing U-Net variant (UNet-CN-CP). The results underscore the importance of evaluating training and inference together. A model that trains cheaply may prove inefficient when repeatedly deployed, while others entail higher upfront cost but amortise more effectively in deployment. Amortising development costs across the model lifecycle is one way to factor in these considerations (Cottier et al., 2024).
FM stands out as an outlier: while competitive in predictive performance, its reliance on numerical integration makes inference substantially more expensive. This is in spite of it operating on downsampled data, as the original resolution was computationally prohibitive (Dao et al., 2023).
Several key points follow from these results. First, model selection cannot be based on efficacy alone, as models with similar predictive accuracy can differ substantially in carbon emissions. Second, once both training and inference costs are accounted for, identifying a single “best” model becomes less straightforward, as illustrated by the incompressible shear flow experiment. A model with a small carbon footprint during training may be unsuitable in deployment, making carbon-aware evaluation across the full lifecycle essential.
The implications extend well beyond our test-bed. Many scientific and engineering problems are governed by well-understood structures, and overlooking them is wasteful. At the same time, tailoring models too closely to the specific physics can limit applicability and can be challenging for complex systems. This motivates the development of general-purpose models such as FNOs and FM, which allow physics priors to be incorporated more easily while retaining flexibility across tasks. However, as demonstrated by FM, alignment with the underlying physics does not guarantee carbon efficiency. Physics should not be incorporated at any cost, but only where it provides clear benefits in terms of predictive performance or carbon efficiency. While our discussion centres on physics-inductive biases, the same logic extends to other forms of prior knowledge: any regularising bias that constrains the learning task appropriately can potentially offer similar trade-offs between efficacy and efficiency.
Ultimately, our results argue for a shift in mindset. Progress in ML should not be equated with ever-larger datasets, models, and compute budgets, but rather evaluated in terms of the value gained relative to the cost incurred. Learning everything from data while ignoring domain knowledge reflects an abundance mindset, in which resources are treated as unlimited. A more sustainable alternative is to exploit what we already know by embedding this knowledge into models. Trading carbon for physics offers a principled path towards architectures that leverage prior knowledge, reduce reliance on brute-force compute, and achieve more favourable efficacy-efficiency trade-offs.
Limitations.
This study has some limitations. First, all models were trained without extensive hyperparameter tuning. This choice emphasises architectural comparison over peak performance, trading statistical robustness for a broader perspective on performance–carbon trade-offs. In practice, however, hyperparameter search can dominate computational cost, and complex models typically require many more training runs during development than simpler baselines. Second, we restricted our study to spatio-temporal forecasting. While this limits the scope of the contribution, the central message about taking efficiency considerations into account by including appropriate inductive biases is applicable across ML tasks. Finally, our measurements do not account for hardware specific optimisations, where certain operations are more efficiently executed on specific accelerators. This, for instance, could yield an advantage to deep learning models compared to PDE solvers.
7 Conclusions
The primary message in this work, backed by empirical evidence, is to increase resource-awareness when designing ML experiments. While this can be accomplished in several ways (Bartoldson et al., 2023), we focus on the role of physics inductive biases and illustrate the different decisions practitioners have to consider. Through multiple experiments from simple oscillators to complex shear flow, we showed how physics-inductive biases affect the trade-off between predictive performance and carbon footprint.
This work advances a perspective grounded in resource-awareness and calls for a shift in mindset. As a community, we should discourage the pursuit of marginal improvements in performance that come from disproportionality higher resource costs. With our work we have made the case to encourage resource-awareness by utilising existing domain knowledge and/or other inductive biases.
Acknowledgements
SW and RS acknowledge funding received from Independent Research Fund Denmark (DFF) under grant agreement number 4307-00143B. JHC acknowledges funding by DFF through the project Arctic Push, grant number 4258-00025B. RS also acknowledges funding received under European Union’s Horizon Europe Research and Innovation Action programme under grant agreements No. 101070284, No. 101070408 and No. 101189771.
The authors thank Erik B Dam, Pedram Bakhtiarifard, Yijun Bian and others from the Machine Learning Section at UCPH for constructive feedback on the manuscript. The authors thank members of SAINTS Lab for useful discussions throughout.
Appendix
Appendix A Flow Matching
We condition on two consecutive states $(u_{0},u_{1}):=(u_{k},u_{k+1})$ such that the path interpolates between current state $u_{0}$ and the target state $u_{1}$. Following Lim et al. (2025), we adopt their interpolation path:
$p_{\tau}(u\mid u_{0},u_{1})=\mathcal{N}\!\left(u\;\middle|\;(1-\tau)u_{0}+\tau u_{1},\;\big(\sigma_{\min}^{2}+\sigma^{2}\tau(1-\tau)\big)I\right),$ (2)
with $\sigma_{\min}=10^{-3}$ and $\sigma=10^{-1}$, as recommended for Navier–Stokes data. The mean interpolates linearly between $u_{0}$ and $u_{1}$, while the variance follows a Brownian-bridge profile (maximal halfway at $\tau=0.5$, minimal at endpoints at $\tau=0$ and $\tau=1$).
The corresponding vector field is:
$v_{\tau}(u\mid u_{0},u_{1})=u_{1}-u_{0}+\frac{\sigma^{2}}{2}\,\frac{1-2\tau}{\sigma_{\min}^{2}+\sigma^{2}\tau(1-\tau)}\Big(u-\big((1-\tau)u_{0}+\tau u_{1}\big)\Big).$ (3)
The parametric vector field $v_{\tau}^{\theta}$ is trained to approximate $v_{\tau}$ by minimizing:
$\mathcal{L}(\theta)=\mathbb{E}_{\tau,\;p_{\tau}(u\mid u_{0},u_{1}),\;q(u_{0},u_{1})}\Big\|v_{\tau}^{\theta}(u)-v_{\tau}(u\mid u_{0},u_{1})\Big\|^{2}.$ (4)
Here $\tau\sim\mathcal{U}[0,1]$, $u\sim p_{\tau}(u\mid u_{0},u_{1})$, and $(u_{0},u_{1})\sim q(u_{0},u_{1})$, where $q$ denotes the empirical distribution of consecutive state pairs in the dataset.
In addition to the conditioning pair $(u_{0},u_{1})$, we provide temporal context by sampling one additional state $u_{c}\in\{u_{k-3},u_{k-2},u_{k-1}\}$, giving a maximum horizon $h=4$ consistent with FNOs and U-nets. All inputs are downsampled by a factor of 16 via average pooling, and predictions are upsampled to the original resolution with bilinear interpolation.
During inference, given a state $u_{k}$, the learned flow is simulated by numerically integrating $v_{\tau}^{\theta}$ from $\tau=0$ to $\tau=1$. We use Euler’s method with $10$ integration steps, which produces intermediate states approximating the continuous trajectory. The final iterate is then upsampled to the original resolution, yielding the forecast $\hat{u}_{k+1}$.
Appendix B Toy Experiment on Harmonic Oscillator
We design a toy experiment to study the effect of physics inductive biases for extrapolation in dynamical systems. The task is to model the trajectory of a harmonic oscillator, comparing standard feedforward networks with variants that incorporate different inductive biases. Specifically, we evaluate four networks with identical architecture but different activation functions, ReLU, Tanh, Sine, and Snake (Ziyin et al., 2020), where:
$\text{Sine}(x)=\sin(x),\qquad\text{Snake}(x)=x+\sin^{2}(x).$ (5)
The Sine and Snake activations introduce oscillatory biases aligned with the dynamics of the oscillator. In addition, we include a PINN that augments the data loss with a PDE residual while using Snake activations.
The governing dynamics are given by the damped harmonic oscillator:
$m\frac{d^{2}u}{dt^{2}}+\mu\frac{du}{dt}+ku=0\penalty 10000\ ,$ (6)
$u(0)=1,\penalty 10000\ \penalty 10000\ \mathrm{and}\penalty 10000\ \penalty 10000\ \frac{du}{dt}=0.$ (7)
The damping ratio is defined as $\delta=\mu/(2m)$. We adopt non-dimensionalized units with $m=1$ and $\omega_{0}=\sqrt{k/m}=1$. We consider both the undamped case ($\delta=0$) and a damped case ($\delta=0.02$).
Training data are sampled from $t\in[-5\pi,5\pi]$ with Gaussian noise ($\mu=0$$\sigma=0.3$). The purely supervised networks use $10,000$ supervised samples, while the PINN uses $100$ points ($1\%$) for the physics loss. Test data are noise-free and sampled over $t\in[5\pi,25\pi]$ to assess extrapolation. The points for evaluating the physics-based loss are sampled across the entire domain.
The purely supervised networks minimize the MSE:
$\mathcal{L}_{\text{NN}}=\mathcal{L}_{\text{data}}=\frac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2},$ (8)
where $u$ denotes the ground truth and $\hat{u}$ the model predictions.
The PINN combines the same data term with a differential equation (DE) residual:
$\mathcal{L}_{\text{PINN}}=\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{pde}}=\frac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2}+\frac{1}{N_{\text{de}}}\sum_{j=1}^{N_{\text{de}}}\left(\frac{d^{2}\hat{u}_{j}}{dt^{2}}+\mu\frac{d\hat{u}_{j}}{dt}+k\hat{u}_{j}\right)^{2}.$ (9)
Further architectural and optimisation details are given in Sec. 4.1.
Appendix C Loss Functions Used in the viscous Burgers’ Equation Experiment
The loss function for the purely data-driven NN, denoted $\mathcal{L}_{\text{NN}}$, is the MSE loss given by:
$\mathcal{L}_{\text{NN}}$ $=\mathcal{L}_{\text{data}}=\dfrac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2}.$ (10)
Here $u$ and $\hat{u}$ denote the ground truth and the predictions.
The U-PINN optimizes solely using the governing equations including those that describe the boundary and initial conditions to ensure a unique solution. Its loss function, $\mathcal{L}_{\text{U-PINN}}$, is composed of the residuals of the PDE and the deviations from the initial and boundary conditions, expressed as:
$\mathcal{L}_{\text{U-PINN}}$ $=\mathcal{L}_{\text{pde}}+\mathcal{L}_{\text{ic}}+\mathcal{L}_{\text{bc}}$ (11)
$=\frac{1}{N_{\text{pde}}}\sum_{i=1}^{N_{\text{pde}}}\left(\frac{\partial\hat{u}_{i}}{\partial t}+\hat{u}_{i}\frac{\partial\hat{u}_{i}}{\partial x}-\nu\frac{\partial^{2}\hat{u}_{i}}{\partial x^{2}}\right)^{2}$ (12)
$+\frac{1}{N_{\text{ic}}}\sum_{j=1}^{N_{\text{ic}}}\left(\hat{u}(x_{j},0)+\sin(\pi x_{j})\right)^{2}$
$+\frac{1}{N_{\text{bc}}}\sum_{k=1}^{N_{\text{bc}}}\left(\hat{u}(-1,t_{k})^{2}+\hat{u}(1,t_{k})^{2}\right).$
The S-PINN combines both data and physics losses. Its total loss function, $\mathcal{L}_{\text{S-PINN}}$ is given by:
$\mathcal{L}_{\text{S-PINN}}=\mathcal{L}_{\text{NN}}+\mathcal{L}_{\text{U-PINN}}=\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{pde}}+\mathcal{L}_{\text{ic}}+\mathcal{L}_{\text{bc}}.$ (13)
Appendix D Incompressible Shear Flow Dataset
This experiment is based on data from The Well (Ohana et al., 2024), consisting of 2D incompressible Navier–Stokes dynamics with an additional passive tracer:
$\partial_{t}\mathbf{u}+\nabla p-\nu\nabla^{2}\mathbf{u}$ $=-\mathbf{u}\cdot\nabla\mathbf{u},$ (14)
$\partial_{t}s-D\nabla^{2}s$ $=-\mathbf{u}\cdot\nabla s,$ (15)
$\nabla\cdot\mathbf{u}$ $=0.$ (16)
Here $\mathbf{u}=(u_{x},u_{y})$ is the velocity components, $p$ is pressure satisfying $\int p=0$, $s$ is the tracer, $\nu$ is viscosity, and $D$ is diffusivity. The PDE parameters are linked to Reynolds and Schmidt numbers as $\nu=1/\text{Re}$ and $D=\nu/\text{Sc}$.
The system is initialized with a horizontal shear flow:
$u_{x}(y)=\tanh\left(5\frac{y-y_{k}}{n_{\text{shear}}w}\right),$ (17)
where $y_{k}$ is the vertical position of the $k$-th shear layer, $n_{\text{shear}}$ is the number of shear layers, and $w$ is a width parameter that controls how sharply the velocity changes across each layer.
To introduce small perturbations in the system, the vertical velocity field $v_{y}$ contains sinusoidal variations along the $x$-direction, localized at the shear layers. These perturbations are given by:
$u_{y}(x,y)=\sin(n_{\text{blobs}}\pi x)\exp\left(\frac{-25}{w^{2}}|y-y_{k}|^{2}\right).$ (18)
The number of oscillations is controlled by $n_{\text{blobs}}$, while $w$ influences how localized these perturbations are around the shear layers.
The tracer field is initialized to match the shear flow, while the pressure field is initialized to zero everywhere in the domain. Thus, the initial conditions are fully characterized by three parameters: $n_{\text{shear}}\in\{2,4\}$, $n_{\text{blobs}}\in\{2,3,4,5\}$, and $w\in\{0.25,0.5,1.0,2.0,4.0\}$. An example of the data is illustrated in Fig. 9.
As part of preprocessing, we standardize each field using $\hat{x}=(x-\mu)/\sigma$ (see Table 2). This normalization balances the contribution of different fields by preventing those with high variance from dominating the loss.
Table 2: Field statistics. Mean and standard deviation of raw fields, averaged over all time steps.
Field Mean Std
s $2.19\times 10^{-3}$ $0.33$
p $1.18\times 10^{-9}$ $0.074$
$\mathbf{v_{x}}$ $2.19\times 10^{-3}$ $0.40$
$\mathbf{v_{y}}$ $1.40\times 10^{-7}$ $0.097$
Figure 9: Example of incompressible shear flow data. Temporal evolution of tracer (s), pressure (p), horizontal velocity ($v_{x}$) and vertical velocity ($v_{y}$) fields at four time steps: $t\in\{0,T/3,2T/3,T\}$. The snapshots illustrate the coupled dynamics of the fields.
Appendix E Flow Matching in Low-Resolution Data Space
Most prior work applies FM in latent space, with a few exceptions (Disch et al., 2025; Baldan et al., 2025) demonstrating its feasibility directly in data space. We adopt this latter strategy to avoid the additional energy and carbon costs of an encoder–decoder pipeline. To make FM computationally tractable, we downsample inputs by a factor of 16, perform training and inference at low resolution, and subsequently upsample to the original resolution. This design inevitably encode a trade-off between efficacy and efficiency: while operating in low-resolution data space suppresses fine-scale dynamics and may introduce artifacts, it delivers substantial efficiency gains, reducing training and inference time by orders of magnitude.
Appendix F Model Configurations
Model configurations are adapted from prior work. U-nets, FNO, and TFNO follow Ohana et al. (2024), UFNO follows Rahman et al. (2023), and FM is adapted from Lim et al. (2025), with some adjustments to match our setting (including operating directly in data space). In line with The Well (Ohana et al., 2024), U-nets and FNO variants are scaled to $\sim$17–20M parameters, while FM is smaller ($\sim$8.5M). Table 3 provides an overview of parameter counts, with detailed configurations summarized in Table 4 (U-nets and FNOs) and Table 5 (FM).
Table 3: Model parameters.
Model #Params
UNet 17.517.5M
UNet-CN 18.618.6M
UNet-CP 17.517.5M
UNet-CN-CP 18.618.6M
FM 8.68.6M
FNO 19.019.0M
TFNO 19.319.3M
UFNO 18.818.8M
Table 4: U-nets and FNOs model configurations.
Model Filter size/modes Init dim Blocks/stage Up/down Bottleneck
UNet, UNet-CP 33 $48$ 11 44 11
UNet-CN, UNet-CN-CP 77 $42$ 22 44 11
FNO, TFNO $16$ $128$ 44
UFNO $8,12,16$ $128$ 11 33 11
Table 5: FM model configuration. The FM model operates in low-resolution data space and is built from transformer stacks with learned positional, timestamp, and distance embeddings.
Model Resolution Init dim Depth Bottleneck Heads
FM $16\times 32$ $256$ 33 44 88
Appendix G Additional Results
This appendix presents supplementary plots that support the findings in the main text. Table 6 reports both the training and inference emissions alongside run times.
Training Inference (one-step)
Model CO2eq (kg) Time (h) CO2eq (g) Time (min)
UNet $2.72$ $12.7$ $3.68\pm 0.03$ $1.0$
UNet-CN $4.71$ $15.9$ $7.77\pm 0.05$ $1.9$
UNet-CP $3.81$ $14.2$ $4.06\pm 0.05$ $1.0$
UNet-CN-CP $5.89$ $20.9$ $4.87\pm 0.02$ $1.2$
FM $3.31$ $14.0$ $23.54\pm 0.13$ $4.6$
FNO $0.80$ $4.0$ $7.87\pm 0.94$ $2.0$
TFNO $1.09$ $5.3$ $7.21\pm 0.12$ $1.8$
UFNO $1.12$ $3.6$ $9.00\pm 0.88$ $2.4$
Table 6: Carbon footprint during training and inference. Reported CO2eq emissions and runtime for all models, measured with CarbonTracker. Inference results correspond to one-step forecasts over the full test set (4,704 samples) and includes the initial model loading cost. Each inference experiment was repeated three times; CO2eq is reported as mean and standard deviation, and runtimes as mean values.
Fig.10-a reports model performance using the variance-scaled root mean squared error (VRMSE), which normalises prediction error by the variance of the ground truth to account for scale differences across fields:
$\text{VRMSE}(u,\hat{u})=\left(\frac{\langle|u-\hat{u}|^{2}\rangle}{\langle|u-\langle u\rangle|^{2}\rangle+\epsilon}\right)^{1/2},$ (19)
where $\langle\cdot\rangle$ denotes spatial averaging, $u$ the ground truth, $\hat{u}$ the prediction, and $\epsilon=10^{-8}$ a small constant for stability. The VRMSE measures errors relative to the natural variability of the data, with $\text{VRMSE}(y,\langle y\rangle)\approx 1$ implying that values greater than one indicate prediction errors larger than typical fluctuations. The results follow the same overall trends as in Fig.8-a, where performance was measured with the Pearson correlation coefficient.
Figure 10-b reports the inference carbon footprint for $24$ autoregressive rollouts of $50$ steps each. Compared to one-step inference over the full test set ($4,704$ samples), the rollout costs tend to average out across models, since the fixed overhead of model loading constitutes a larger share of the total cost. Flow matching remains an outlier, standing out as the most expensive model during inference.
Figure 10: Predictive performance and rollout carbon footprint. a: VRMSE vs. training carbon footprint. Each point shows every second step of a 20-step rollout, averaged across the four fields (lighter shades indicate earlier steps). b: Carbon footprint of inference during rollouts. Reported CO2eq emissions for 24 autoregressive rollouts of 50 steps. Inference costs include the initial model loading cost. Models are ordered from no bias (left) to strong bias (right).
Figs. 11, 12, and 13 provide rollout prediction examples for the three physical fields not shown in the main text (pressure fields and the two velocity components). The observations made for the tracer field in the main text extend to these additional fields.
Figure 11: Example of rollout predictions for pressure field. Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout. Models labels are ordered from no bias (top) to strong bias (bottom).
Figure 12: Example of rollout predictions for horizontal velocity field. Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout. Models labels are ordered from no bias (top) to strong bias (bottom).
Figure 13: Example of rollout predictions for vertical velocity field. Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout. Models labels are ordered from no bias (top) to strong bias (bottom).

### v2

@@
-Trading Carbon for Physics: On the Resource Efficiency of Machine Learning for Spatio-Temporal Forecasting
+Physics Priors Offer Useful Accuracy-Carbon Trade-Offs in Spatio-Temporal Forecasting
 Abstract
 Development of modern deep learning methods has been driven primarily by the push for improving model efficacy (accuracy metrics).
-This sole focus on efficacy has steered development of large-scale models that require massive resources, and results in considerable carbon footprint across the model life-cycle.
+This sole focus on efficacy has steered development of large-scale models that require massive computational resources, and results in considerable energy consumption and corresponding carbon footprint across the model lifecycle.
 In this work, we explore how physics inductive biases can offer useful trade-offs between model efficacy and model efficiency (compute, energy, and carbon).
-We study a variety of models for spatio-temporal forecasting, a task governed by physical laws and well-suited for exploring different levels of physics inductive bias.
-We show that embedding physics inductive biases into the model design can yield substantial efficiency gains while retaining or even improving efficacy for the tasks under consideration.
-In addition to using standard physics-informed spatio-temporal models, we demonstrate the usefulness of more recent models like flow matching as a general purpose method for spatio-temporal forecasting.
-Our experiments show that incorporating physics inductive biases offer a principled way to improve the efficiency and reduce the carbon footprint of machine learning models.
+We study models with strong, weak, and no physics-inductive biases for spatio-temporal forecasting of incompressible shear flow, a task governed by the Navier-Stokes equations.
+We find that models with stronger physics priors achieve substantially lower training footprints, but this advantage does not straightforwardly extend to inference, highlighting the importance of evaluating carbon costs across the full model lifecycle rather than any single stage.
 We argue that model efficiency, along with model efficacy, should become a core consideration driving machine learning model development and deployment.
-[^1] Source code for the experiments in this work are available at: https://github.com/sophiawilson18/FlowMatching.
 1 Introduction
-Machine learning (ML) has advanced rapidly over the past decade (Sevilla et al., 2022), driven largely by the pursuit of improving model efficacy (accuracy metrics).
-This narrow focus has led to increasingly large models that require substantial computational resources and generate significant carbon emissions across their lifecycle (Strubell et al., 2019; Anthony et al., 2020; Luccioni et al., 2023).
+Machine learning (ML) has advanced rapidly over the past decade, driven largely by the pursuit of improving model efficacy (accuracy metrics).
+This narrow focus has led to increasingly large models that require substantial computational resources and generate significant carbon emissions across their lifecycle [1, 2, 3, 4].
 As models scale, efficacy gains diminish while costs escalate, underscoring the need to move beyond accuracy as the sole measure.
-A more holistic assessment must balance efficacy with efficiency, capturing not only predictive performance but also compute, energy, and carbon costs (Henderson et al., 2022).
-This requires exploring trade-offs and identifying strategies that strike a favourable balance, rather than lopsided optimization.
+A more holistic assessment must balance efficacy with efficiency, capturing not only predictive performance but also compute, energy, and carbon costs [5].
+This requires identifying strategies that strike a favourable balance rather than optimising for accuracy alone.
 Figure 1: a: Conceptual map of modelling approaches by data availability (vertical axis) and domain knowledge (horizontal axis).
 Physics-informed ML lies in the middle, combining data-driven flexibility with physics priors.
-b: Test MSE versus CO2eq emissions for a data-driven NN (red squares), an unsupervised PINN (blue dots), and a semi-supervised PINN (grey triangles).
+b: Test mean squared error (MSE) versus CO2eq emissions for a data-driven NN (red squares), an unsupervised PINN (blue dots), and a semi-supervised PINN (grey triangles).
 The dashed line marks the Pareto front, dominated by the S-PINN, indicating the best accuracy-carbon trade-off.
-Details of the experiment are in Sec.
-4.2.
-Physics-informed ML offers a promising way to achieve this balance.
-Physics priors such as conservations laws, symmetries, or governing equations can be embedded into model design through, e.g., loss terms, architectural layers, or mechanisms.
-This combines the flexibility of data-driven methods with the structure of domain knowledge (Fig.1-a).
-As shown in Fig.
-1-b, a semi-supervised physics-informed neural network (PINN) achieves the best accuracy–carbon trade-off, outperforming both a purely data-driven neural network (NN) and a physics-only PINN (see Sec.
-4.1 for details).
-This example highlights how hybrid approaches can yield superior efficiency without sacrificing efficacy.
+Details of the experiment are in App.
+0.A.
+Physics-informed ML offers a promising path towards this balance.
+By embedding prior knowledge about the underlying system directly into the model, one can combine the flexibility of data-driven methods with the structure of domain knowledge, as illustrated in Fig.
+1-a.
+Prior work has shown that embedding conservation laws, symmetries, or governing equations can reduce the number of trainable parameters [6, 7], reduce the amount of training data [8, 9, 10], and accelerate training convergence [11, 12].
+Whether physics-informed ML models can also reduce energy consumption and carbon footprint remains a less explored area.
+Fig.
+1-b offers a concrete illustration of its potential.
+In a simple experiment on a partial differential equation (PDE) system, we compare three models that share the same underlying architecture but differ in how much they rely on physics through their loss functions: a purely data-driven neural network (NN), an unsupervised physics-informed neural network (PINN) trained solely on PDE residuals, and a semi-supervised PINN combining data and physics.
+The semi-supervised PINN dominates the Pareto front, achieving lower prediction error at a lower carbon cost than either alternative.
+This suggests that striking the right balance between data and physics does not just improve accuracy, but can simultaneously reduce the carbon footprint.
+Full details of this experiment are provided in App.
+0.A.
 Building on this intuition, we extend the perspective from task-specific models such as PINNs to general-purpose architectures that embed physics inductive biases in ways that are transferable across datasets and domains.
-We focus on spatio-temporal forecasting of partial differential equations (PDE) dynamics, a setting characterised by high computational cost and practical relevance.
-Operational weather and climate forecasting, for instance, requires large-scale PDE solvers running on supercomputers multiple times per day (Lam et al., 2023).
-In such contexts, even modest efficiency improvements can translate into substantial operational carbon savings, making this domain an ideal test-bed for exploring sustainable ML design (Van Wynsberghe, 2021).
+We focus on spatio-temporal forecasting of PDE dynamics, a setting characterised by high computational cost and practical relevance.
+Operational weather and climate forecasting, for instance, requires large-scale PDE solvers running on supercomputers multiple times per day [13].
+In such contexts, even modest efficiency improvements can translate into substantial operational carbon savings [14], making this domain an ideal test-bed for exploring sustainable ML design [15].
 Our key contributions towards this end are:
   1. 1.
-    Carbon-aware evaluation: We promote evaluation practices that consider both predictive accuracy and carbon cost, explicitly characterising trade-offs.
+    Carbon-aware evaluation: We promote evaluation practices that explicitly account for training and inference costs alongside predictive accuracy.
   2. 2.
-    Characterising the accuracy–carbon trade-off: We conduct three experiments on PDE dynamics and spatio-temporal forecasting to compare a range of models with varying levels of physics inductive bias.
-Our main study focuses on forecasting incompressible shear flow, comparing U-nets (no/weak bias), flow matching (medium bias), and Fourier neural operators (strong bias).
-Figure 2: Modelling the harmonic oscillator. Training and extrapolation predictions when using NNs with different activation functions: ReLU, Tanh, Sine, Snake, and using a PINN.
+    Characterising the accuracy-carbon trade-off: We provide an empirical characterisation of how varying degrees of physics priors affect the accuracy-carbon trade-off in spatio-temporal forecasting, evaluated on a complex, high-resolution dataset of incompressible shear flow.
 2 Background and Related Work
+Figure 2: Bias-variance trade-off.
+Top: classical trade-off with minimum total error at optimal model complexity.
+Bottom: possible analogous balance between physics-based constraints and data-driven flexibility.
 Revisiting the Bias-Variance Trade-Off.
-It is commonly understood in ML that strong model assumptions (bias) risk under-fitting, whereas high sensitivity to training data can lead to over-fitting (variance) (Kohavi et al., 1996).
+It is commonly understood in ML that strong model assumptions (bias) risk under-fitting, whereas high sensitivity to training data can lead to over-fitting (variance) [16].
 Striking the right balance is crucial to obtain the right class of models.
-The same trade-off can also offer a useful way to perform model selection when viewed through the lens of resource consumption.
-Typically, if domain knowledge (physics-based or others) can be incorporated into models as inductive bias, it can reduce the reliance on training data or reduce prediction errors.
-This is illustrated in Fig.
-2 for the task of modelling a harmonic oscillator.
-Models with appropriate inductive bias (sinusoidal oscillations) perform better for the same number of training data points and model parameters.
-See Sec.
-4.1 for more details.
+We draw an analogy to model design under resource constraints.
+Incorporating domain knowledge as inductive bias can reduce reliance on large datasets and potentially lower computational demands, but may limit model flexibility.
+Conversely, highly data-driven models offer flexibility, but must learn all patterns from data which often requires lots of data and computational resources.
+We therefore hypothesise that a physics-data balance may exist, analogous to the bias-variance trade-off (Fig.
+2).
 The Question of Resource-Awareness.
 The standard metrics used to characterise the performance of ML models primarily focus on efficacy and are untethered from efficiency considerations.
-This introduces a preference for models that could offer high efficacy at large resource costs (Bakhtiarifard et al., 2024).
+This introduces a preference for models that could offer high efficacy at large resource costs [17].
 There are two primary streams of work that aim to include resource-aware metrics: using composite metrics and multi-objective optimisation.
@@
 Evchenko et al.
-(2021) introduced a composite metric that combines efficacy and efficiency metrics to study the influence of resource constraints on traditional (non deep learning) ML models.
-Carburacy (Moro et al., 2023) extends the idea to transformer models by jointly quantifying accuracy and carbon emissions.
+[18] introduced a composite metric that combines efficacy and efficiency metrics to study the influence of resource constraints on traditional (non deep learning) ML models.
+Carburacy [19] extends the idea to transformer models by jointly quantifying accuracy and carbon emissions.
 More recently, Kapoor et al.
-(2025) proposed EcoL2, a composite metric similar to Carburacy but extended to include the costs of data generation and hyperparameter tuning.
-Multi-objective optimisation has been primarily used in neural architecture search (NAS) resulting in models like EfficientNet (Tan and Le, 2019).
-More recently, energy consumption (Bakhtiarifard et al., 2024) and carbon footprint (Zhao et al., 2024) have been taken into consideration in NAS.
+[20] proposed EcoL2, a composite metric similar to Carburacy but extended to include the costs of data generation and hyperparameter tuning.
+Multi-objective optimisation has been primarily used in neural architecture search (NAS) resulting in models like EfficientNet [21].
+More recently, energy consumption [17] and carbon footprint [22] have been taken into consideration in NAS.
 These methods, however, have not actively studied the influence of physics inductive biases on the resource consumption of ML models.
 Spatio-Temporal Forecasting Models.
-There are many model classes for the task of spatio-temporal forecasting; each differing in the extent of physics inductive bias they use.
-Fig.
-3 provides a high-level overview of this span for models considered in this work.
-Figure 3: Spectrum of model families from purely data-driven ML to physics-based solvers.
-U-net.
-U-nets (Ronneberger et al., 2015) can be used for spatio-temporal forecasting by mapping a window of past states to future ones.
-They capture local correlations by fusing features across multiple spatial scales, but do not explicitly constrain the dynamics beyond what is enforced by the training data.
-Mild forms of inductive biases can be introduced to convolutional architectures like U-net by including circular padding to respect periodic boundaries and avoid edge discontinuities.
-Flow Matching.
-Recent classes of probabilistic generative models like flow matching (FM) (Lipman et al., 2023) can be viewed as an intermediate class of models that can include some degree of inductive bias by modelling dynamics as continuous-time flows.
-This formulation resonates with dynamical systems and intuitively suggests smoother trajectories.
-FM has been applied to spatio-temporal forecasting and other PDE dynamics tasks (Lim et al., 2025; Li et al., 2025), with Baldan et al.
-(2025) further embedding PDE residuals into the objective.
-Fourier Neural Operator.
-The Fourier Neural Operator (FNO) (Kovachki et al., 2023) introduces a physics prior through its Fourier-space representation, enabling the model to capture global spectral interactions.
+There are many model classes for the task of spatio-temporal forecasting, differing in the extent of physics inductive bias they use.
+In this work, we focus on two: the U-net [23] and the Fourier Neural Operator (FNO) [24].
+U-nets perform forecasting by mapping a window of past states to future states.
+They capture local correlations by fusing features across multiple spatial scales, but do not impose any explicit constraints on the underlying dynamics beyond those learning from data.
+Mild forms of inductive biases can be introduced to convolutional architectures like U-net, for example through circular padding to respect periodic boundaries and avoid edge discontinuities.
+The FNO incorporates a stronger inductive bias through its Fourier-space representation, enabling the model to capture global spectral interactions.
 This makes it particularly well-suited to systems with periodicity and long-range correlations.
-In our spatio-temporal forecasting setting, we adapt the FNO from its original operator-learning formulation (Lu et al., 2021) to an autoregressive setup, where its spectral representation provides a strong inductive bias compared to the U-net by explicitly encoding global and periodic structure.
-Physics-Informed Neural Network.
-PINNs (Raissi et al., 2019) embed governing PDEs directly into the training objective, thereby regularising the optimisation landscape and guiding solutions towards physical consistency.
-By minimizing PDE residuals, PINNs can even be trained without supervised data, representing the strongest inductive bias among the models considered.
-Embedding conservation laws, symmetries, or governing PDEs into the ML model development pipeline has been shown to reduce the number of trainable parameters (Dutta et al., ; Patra et al., 2024), reduce the amount of training data (Psaros et al., 2023; Zhong et al., 2021), and accelerate convergence (Brehmer et al., 2024; Jahani-nasab and Bijarchi, 2024).
-3 Methods
+In our setting, we adapt the FNO from its original operator-learning formulation [25] to an autoregressive setup, where its spectral parameterisation explicitly encodes global and periodic structure, in contrast to the primarily local bias of the U-net.
+3 Experiment
 Problem Setup.
-We consider spatio-temporal forecasting of dynamical systems in the general setting, where the objective is to approximate the solution $u(x,y,t)$ of a PDE given basic inputs such as initial and boundary conditions.
+We consider spatio-temporal forecasting of dynamical systems, where the objective is to approximate the solution $u(x,y,t)$ of a PDE given initial and boundary conditions.
 We frame this as an autoregressive prediction task on a two-dimensional, uniformly spaced mesh $(x,y)\in\{x_{1},x_{2},\ldots,x_{H}\}\times\{y_{1},y_{2},\ldots,y_{W}\}$ with discretized time steps $t\in\{t_{1},t_{2},\ldots,t_{T}\}$.
-At each step $t_{k}$ the state is represented as a tensor $u_{k}(x,y)\in\mathbb{R}^{H\times W\times C}$, where $k\in\{1,2,..,T\}$ indexes discrete time steps in a trajectory of length $T$.
+At each step $t_{k}$ the state is represented as a tensor $u_{k}(x,y)\in\mathrm{R}^{H\times W\times C}$, where $k\in\{1,2,..,T\}$ indexes discrete time steps in a trajectory of length $T$.
 $H$ and $W$ are spatial dimensions and $C$ is the number of physical fields (channels), such as pressure or velocity components.
 To capture temporal dependencies, the model is conditioned on a finite history of length $h$.
-The input window is denoted as $u_{k-h+1:k}(x,y):=(u_{k-h+1}(x,y),\ldots,u_{k}(x.y))$.
+The input window is denoted as $u_{k-h+1:k}(x,y):=(u_{k-h+1}(x,y),\\ \ldots,u_{k}(x.y))$.
 A parameterised function $f_{\theta}$ is trained to predict:
@@
 1 and feeding the model’s own predictions back as inputs.
-We adapt this generalised model formulation to three problems.
-Each problem is then tackled with relevant set of methods that range from being entirely data-driven to offering relevant degrees of physics inductive biases.
-  1. 1.
-    Harmonic Oscillator.
-Modelling the dynamics of the harmonic oscillator only requires the time variable i.e, $u(t)$.
-For the harmonic oscillator, Eq.
-1 reduces to $u_{k+1}=f_{\theta}(u_{k-h+1:k}\big)$, where the spatial variables $(x,y)$ are dropped.
-    We assess the impact of appropriate physics-inductive biases by comparing simple feed-forward NNs with and without biases.
-Specifically, these biases are introduced through activation functions and a loss term, for modelling the damped and undamped oscillators.
-  2. 2.
-    Viscous Burgers’ Equation.
-Solving the viscous Burgers’ equation increases the modelling complexity compared to the harmonic oscillator, as it has one spatial and one temporal variable to be considered.
-Eq.
-1 for this case is given as: $u_{k+1}(x)=f_{\theta}(u_{k-h+1:k}(x)\big)$, where only one spatial variable $x$ is modelled along time.
-    We evaluate how different combinations of data and physics constraints influence predictive accuracy and efficiency by comparing a supervised NN, a physics-only PINN, and a hybrid PINN for approximating the viscous Burgers’ equation.
-  3. 3.
-    Incompressible Shear Flow.
-Modelling the spatio-temporal dynamics of incompressible shear flow governed by Navier-Stokes equations is the most complex task considered.
-It has dependence on both the spatial and temporal variables.
-As a result, we use Eq.
-1 in its original formulation.
-    We evaluate several models spanning a broad spectrum of physics biases, depicted in Fig.
+We instantiate this framework for forecasting incompressible shear flow governed by the Navier-Stokes equations, a setting that is both computationally demanding and governed by well-understood physical structure, making it well-suited for studying the efficacy-efficiency trade-off.
+Data.
+We use the incompressible shear flow data from The Well collection [26], generated with Dedalus [27].
+The dataset captures the coupled non-linear evolution of four physical fields: horizontal and vertical velocity, pressure, and a passive tracer on a periodic spatial domain, see example in Fig.
 3.
-Purely data-driven models like U-net, the moderately biased model FM, and strongly biased neural operators are adapted for this task.
-The U-nets and FNOs use a history of $h=4$ states, i.e.
-they are trained for one-step prediction according to Eq.
-1 i.e., $u_{k+1}(x,y)=f_{\theta}(u_{k-3:k}(x,y)\big).$
-    FM replaces direct one-step prediction with a continuous transport between consecutive states.
-Instead of learning a map $u_{k-h+1:k}(x,y)\mapsto u_{k+1}(x,y)$, FM predicts the next state by transporting the current state forward along a learned flow.
-Starting from $u_{k}(x,y)$, the method defines a flow trajectory $\phi_{\tau}(u_{k})$ over a synthetic time axis $\tau\in[0,1]$, where the flow evolves under a time-dependent vector field $v_{\tau}$.
-The one-step forecast is then obtained as the endpoint of this flow: $\hat{u}_{k+1}(x,y)=\phi_{1}(u_{k}(x,y))$, where $\partial_{\tau}\phi_{\tau}(u)=v_{\tau}(\phi_{\tau}(u)),\phi_{0}(u)=u_{k}(x,y).$ In practice, a NN is trained to approximate this vector field, defined by an interpolation path between consecutive states.
-Further FM details including the interpolation path, vector field, and learning objective are provided in App.
-A.
-4 Data and Experiments
-Based on the problem formulation in Sec.
-3, we set-up experiments on the three tasks to explore the trade-off between efficacy and efficiency.
-Specific choices for each of the experiments are elaborated in detail next.
-4.1 Harmonic Oscillator
-We use simple feed-forward NNs to model the harmonic oscillator by introducing physics biases in two ways: through activation functions, where Sine and Snake (Ziyin et al., 2020) have periodic biases, and through a physics-informed loss objective.
-ReLU and Tanh serve as unbiased baselines.
-The training data is generated from the damped harmonic oscillator, $u(t)=e^{-\delta t}(2A\cos(\phi+\omega t))$, where: $A=1/(2\cos\phi),\penalty 10000\ \penalty 10000\ \phi=\arctan\left(-\delta/\omega\right)\penalty 10000\ \penalty 10000\ \mathrm{and}\penalty 10000\ \penalty 10000\ \omega=\sqrt{\omega_{0}^{2}-\delta^{2}}\penalty 10000\$, with added Gaussian noise $(\mu=0,\sigma=0.3)$.
-We consider both the non-damped ($\delta=0$) and the damped ($\delta=0.02$) regimes.
-All models use the same underlying architecture (two hidden layers of width 64), trained for $100$ epochs with Adam optimizer (Kingma and Ba, 2015) using a batch size of 25.
-Learning rates are $0.01$ for standard networks and $0.001$ for the PINN, which employs Snake activations.
-Results are averaged over $10$ runs.
-Additional details are provided in App.
-B.
-4.2 Viscous Burgers’ Equation
-Figure 4: Target and training data for viscous Burgers’ equation. a: Target data. b: Supervised samples. c: PDE residuals samples. d: Initial and boundary condition points.
-In this experiment, we study models to approximate the solution of the viscous Burgers’ equation with Dirichlet boundary conditions on a 2D spatio-temporal grid:
-$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}=0,$
-$u(x,0)=-\sin(\pi x),$
-$u(1,t)=u(-1,t)=0.$
-for $x\in[-1,1]$, $t\in[0,1]$, and with fixed viscosity ($\nu=0.025$).
-The domain is discretized with $256$ spatial points and $100$ time steps.
-We compare three models with varying physics inductive biases, all using the same underlying architecture (comprising four hidden layers, width $50$, tanh activation functions):
-  • NN: a supervised NN trained with the mean-squared error (MSE) on ground-truth data, using the loss objective: $\mathcal{L}_{\text{NN}}=\mathcal{L}_{\text{data}}$.
-  • U-PINN: an unsupervised PINN trained only on physics constraints that minimizes PDE residuals, initial conditions (IC) and boundary conditions (BC).
-Its loss function, $\mathcal{L}_{\text{U-PINN}}$, is composed of the residuals of the PDE and the deviations from the initial/boundary conditions, expressed as: $\mathcal{L}_{\text{U-PINN}}=\mathcal{L}_{\text{PDE}}+\mathcal{L}_{\text{IC}}+\mathcal{L}_{\text{BC}}$.
-  • S-PINN: a semi-supervised PINN where half of the training points are used for PDE residuals and boundary/initial conditions, while the other half enforce a data-driven penalty (MSE).
-Its loss objective is given as, $\mathcal{L}_{\text{S-PINN}}=\mathcal{L}_{\text{U-PINN}}+\mathcal{L}_{\text{NN}}$.
-The specific formulations of all three loss objectives are provided in App.
-C.
-All models were trained using Adam optimizer followed by L-BFGS optimizer (Liu and Nocedal, 1989) with learning rate $10^{-3}$.
-Dataset sizes range from $10^{3}$ to $5\times 10^{3}$ training points for PINNs and up to $8\times 10^{3}$ for the NN.
-Data and code are adapted from Khoa (2022).
-4.3 Incompressible Shear Flow
-We set-up the large-scale experiment on modelling incompressible shear flow governed by the Navier–Stokes equations, using the spatio-temporal forecasting models introduced in Section 2.
-Data.
-The incompreesible shear flow data is derived from The Well dataset (Ohana et al., 2024) generated with Dedalus (Burns et al., 2020).
-We use a subset of the full dataset consisting of 240 simulations spanning six PDE parameter settings with $40$ initial conditions each.
+We work with a subset, which consists of 240 simulations spanning six PDE parameter settings and $40$ initial conditions settings.
 Simulations are resolved at $256\times 512$ with $200$ time steps ($\Delta t=0.1$).
-The data is split by initial condition into 80/10/10 for training, validation, and testing, yielding $37,632$ one-step samples for training, $4,704$ for one-step inference, and $24$ trajectories for autoregressive rollout (initialised at $t=0$).
-The dataset is particularly challenging due to the coupled evolution of four physical fields (velocity components, pressure, tracer) and its non-linear spatiotemporal dynamics at high resolution.
-Further details on PDE formulation and initialization are provided in App.
-D.
-Model(s) Physics inductive bias Strength
+The data is split by initial condition into 8080/1010/$10$ for training, validation, and testing, yielding $37,632$ one-step samples for training, $4,704$ for one-step inference, and $24$ trajectories for autoregressive rollout initialised at $t=0$.
+Further details on the PDE formulation and initialisation are provided in App.
+0.B.
+Figure 3: Example of incompressible shear flow data.
+Temporal evolution of tracer ($s$), pressure ($p$), horizontal velocity ($v_{x}$) and vertical velocity ($v_{y}$) fields at four time steps: $t\in\{0,T/3,2T/3,T\}$.
+The snapshots illustrate the coupled dynamics of the fields.
+Table 1: Overview of the seven models and the relative strength of their physics inductive biases.
+Models Physics inductive bias Strength
 UNet, UNet-CN None None
 UNet-CP, UNet-CN-CP Circular padding Weak
-FM Continuous-time vector flow Medium
 FNO, TFNO, UFNO Spectral layers Strong
-Table 1: Overview of the eight models and the relative strength of their physics inductive biases.
 Models.
-We evaluate eight architectures spanning a spectrum of physics inductive biases, from purely data-driven to strongly physics-informed, summarised in Table 1.
-We use standard U-nets (Ronneberger et al., 2015) and a ConvNeXt variant (UNet-CN) (Liu et al., 2022) for improved hardware-efficiency.
-Zero-padding serves as our no-bias baseline, while circular padding (CP) introduces a weak prior consistent with periodic boundary conditions in the dataset.
-We also include our proposed FM model.
-Finally, we include three variants of the Fourier Neural Operator (FNO): a baseline model, a Tucker-factorized (TFNO) variant (Kossaifi et al., 2023) that improves parameter effiency, and a U-shaped (UFNO) variant (Rahman et al., 2023) designed to better capture multi-scale interactions.
+We evaluate seven models spanning a spectrum of physics inductive biases, summarised in Table 1.
+We consider a baseline U-net and a ConvNeXt variant (UNet-CN) [28], which improves hardware efficiency through modern convolutional design.
+We introduce a weak form of inductive bias by applying circular padding to both architectures (UNet-CP, UNet-CN-CP).
+For models with stronger physics priors, we include three FNO variants: a baseline FNO, a Tucker-factorized variant (TFNO) [29] that improves parameter efficiency, and a U-shaped variant (UFNO) [30] designed to better capture multi-scale interactions.
 Training Procedure.
-Across all models, we use a batch size of 16 and the AdamW optimizer (Loshchilov and Hutter, 2019) with $10^{-4}$ weight decay.
-For U-nets and FNOs, we adopt the coarse-tuned learning rates reported in The Well ($5\cdot 10^{-4}$ and $10^{-3}$, respectively).
-For FM, we performed a separate coarse tuning and selected $5\cdot 10^{-3}$.
-A linear warm-up cosine scheduler was applied for the first three epochs.
-The loss optimised is MSE averaged across fields and space.
+All models use a history of $h=4$ states and are trained for one-step prediction according to Eq.
+1.
+Across all models, we use a batch size of 16 and the AdamW optimizer [31] with $10^{-4}$ weight decay.
+We adopt the coarse-tuned learning rates reported in The Well [26] ($5\cdot 10^{-4}$ for U-nets and $10^{-3}$ for FNOs, respectively), which were tuned across several datasets including the one used here.
+A linear warm-up cosine scheduler was applied for the first three epochs and the loss is the mean squared error (MSE) averaged across fields and space.
 Training ran on a single Nvidia A40 GPU ($\leq$ 24h) with early stopping (patience 6).
-FM was trained in low-resolution data space with subsequent upsampling, avoiding the additional overhead of an autoencoder.
-Further details on FM in low-resolution data space are provided in App.
-E, and complete model configurations are listed in App.
-F.
-4.4 Energy and Carbon Metrics
-We employ Carbontracker (Anthony et al., 2020) to estimate the energy consumption and carbon footprint during training and inference by monitoring real-time hardware power usage (CPU, GPU, DRAM) and adjusting for infrastructure overhead via the power usage effectiveness.
-The carbon footprint is reported as carbon dioxide equivalent (CO2eq) emissions, which converts all greenhouse gases into an equivalent CO2 warming potential.
-It is derived from the total energy consumed and by using the global average carbon intensity for 2024 at $445$ gCO2/kWh, ensuring fair comparisons across models (IEA, 2025).
-Figure 5: Damped harmonic oscillator predictions.
-Mean predictions (solid lines) with standard deviation (shaded areas) from $10$ runs.
-Results are shown for the two best-performing models: Snake and PINN.
+Full model configurations are listed in App.
+0.C.
+Carbon Footprint.
+To estimate energy consumption during training and inference, we employ Carbontracker [2].
+We report results in terms of carbon footprint rather than energy consumption directly, since energy is only problematic to the extent that it is generated from carbon-emitting sources.
+Were the grid entirely renewable, the same energy expenditure would carry negligible direct carbon emissions.
+In practice, however, the global energy mix remains far from carbon-neutral, and we therefore convert energy to carbon dioxide equivalent (CO2eq) emissions using the global average carbon intensity for 2024 at $445$ gCO2/kWh [32].
+Using a fixed carbon intensity ensures fair comparisons across models, independent of when or where they were trained.
 4 Results
-5.1 Harmonic Oscillator Toy Experiment
-The results in Fig.
-2 show that models with non-periodic activations perform poorly outside the training domain (extrapolation).
-Periodic activations (Sine, Snake) offer modest improvements, with the model using Snake activation function performing best.
-The PINN achieved superior results in both interpolation and extrapolation tasks; most notably for the damped harmonic oscillator shown in Fig 5.
-It remained stable and accurate even as prediction range increased and dynamics grew more complex.
-5.2 Viscous Burgers’ Equation Experiment
-For the viscous Burgers’ equation experiment, the comparison with NN, U-PINN, and S-PINN show interesting trends.
-Both PINNs achieve consistently lower test losses at lower carbon footprints compared to the supervised NN (see Fig.
-1-b).
-Although the NN has the lowest carbon footprint for fixed epochs and data size, achieving a given accuracy is up to $10$ times more carbon-efficient with the PINNs.
-This S-PINN, which mixes data and physics, dominates the Pareto front in Fig.
-1-b, suggesting that combining physics and data gives the model which balances efficacy and efficiency.
-5.3 Incompressible Shear Flow Experiment
-The first two simple experiments discussed above establish key insights into how physics can affect the model performance and resource consumption for specific tasks.
-In the incompressible shear flow experiment, we extend this analysis to a more complex task using a wide spectrum of models.
+Figure 4: Pearson r over 20 rollout steps. Values are averaged over trajectories and fields. Models labels are ordered from no bias (top) to strong bias (bottom).
 Predictive Performance.
-We first assess predictive performance using Pearson correlation $r$ at each step of a 20-step rollout, see Fig.
-6.
-All eight models (see Table 1) achieve high accuracy for the first few steps (r $\approx$ 1), after which performance declines at different rates.
-Over longer horizons, UNet-CN, UNet-CN-CP, and UFNO emerge as the strongest performers.
-Figure 6: Pearson r over 20 rollout steps. Values are averaged over the four fields: velocity components, tracer, and pressure. Models labels are ordered from no bias (top) to strong bias (bottom).
-Figure 7 illustrates these stability differences visually.
-UNet and UNet-CP deteriorate rapidly, accumulating distortions within just a few steps.
-The modernised variants (UNet-CN, UNet-CN-CP) show marked improvements: UNet-CN captures fine-scale detail but gradually disrupts large-scale dynamics, while UNet-CN-CP better preserves both large and fine-scale structure at the cost of high-frequency artifacts.
-FM exhibits good temporal stability, though it tends to exaggerate some fine-scale structures.
-FNO and TFNO maintain the dominant large-scale dynamics but exhibit a systematic drift toward higher values, fail to capture fine-scale variability and eventually introduce distortions.
-UFNO better recovers both large- and small-scale structures, though these gradually smooth out over time.
-Figure 7: Example rollout predictions for one of the four fields (tracer).
-Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout, illustrating differences in long-horizon stability and fidelity.
-Models labels are ordered from no bias (top) to strong bias (bottom).
-Results for the three other fields are provided in Appendix G.
-Training Cost.
 […]
-Inference Cost.
-To provide a more complete evaluation, we also account for inference costs of all the trained models.
-This can be useful in choosing spatio-temporal forecasting models as inference costs can dominate over training or development costs.
-Figure 8: Predictive performance and carbon footprint.
-a: Predictive accuracy (Pearson r) versus training emissions (kgCO2eq); each point represents every second step of a 20-step rollout (lighter shades indicate earlier steps), averaged over four fields.
-b: Training emissions (kgCO2eq) from a single run per model.
+Carbon Footprint.
+We next evaluate the carbon footprint and runtime associated with both training and inference (Table 2).
+Considering carbon efficiency (CO2eq per epoch), FNO and TFNO are roughly as carbon-intensive as the baseline UNet, while the other U-net variants are considerably more intensive.
+Within the U-net family, both the ConvNext design and circular padding make the model more carbon-intensive, with the exception of UNet-CN-CP, which is more efficient relative to UNet-CN.
+This indicates that architectural design choices can interact with computational efficiency in non-trivial ways.
+UFNO stands out as the by far most carbon-efficient model.
+Table 2: Carbon footprint during training and inference.
+Reported CO2eq emissions and runtime for all models, measured with CarbonTracker.
+Inference results correspond to one-step forecasts over the full test set (4,704 samples) and include the initial model loading cost.
+Each inference experiment was repeated three times; CO2eq is reported as mean and standard deviation, and runtimes as mean values.
+Training Inference (one-step)
+Model CO2eq (kg) Time (h) # Epochs CO2eq/Epoch (g) CO2eq (g) Time (min)
+UNet $2.72$ $12.7$ $43$ $63$ $3.68\pm 0.03$ $1.0$
+UNet-CN $4.71$ $15.9$ $27$ $174$ $7.77\pm 0.05$ $1.9$
+UNet-CP $3.81$ $14.2$ $35$ $109$ $4.06\pm 0.05$ $1.0$
+UNet-CN-CP $5.89$ $20.9$ $45$ $131$ $4.87\pm 0.02$ $1.2$
+FNO $0.80$ $4.0$ $12$ $67$ $7.87\pm 0.94$ $2.0$
+TFNO $1.09$ $5.3$ $16$ $68$ $7.21\pm 0.12$ $1.8$
+UFNO $1.12$ $3.6$ $30$ $37$ $9.00\pm 0.88$ $2.4$
+Figure 5: a: Training emissions (kgCO2eq) from a single run per model.
+b: One-step inference emissions (gCO2eq) from three repeated runs over the full test set; error bars indicate variability across repetitions.
+Note that inference costs also include the initial model loading cost.
+c: Carbon footprint of inference during rollouts.
+Reported CO2eq emissions for 24 autoregressive rollouts of 50 steps.
+Inference costs include the initial model loading cost.
 Models are ordered from no bias (left) to strong bias (right).
-c: One-step inference emissions (gCO2eq) from three repeated runs over the full test set; error bars indicate variability across repetitions.
-Note that inference costs also include the initial model loading cost.
-[…]
-Padding choices further illustrate the sensitivity of emissions to architectural design.
-Switching from zero to circular padding increases training emissions for both UNet and UNet-CN.
-At inference, the effects diverge: circular padding slightly increases the footprint for UNet, while significantly reducing it for UNet-CN.
-This highlights how even minor inductive design choices – in this case enforcing circular padding to reflect periodic boundaries – can shift the efficacy–efficiency balance.
-Table 6 in App.
-G reports both training and inference footprints alongside run times.
-While runtime and energy consumption/ carbon emissions are closely linked, they are not strictly proportional, as some operations could take longer but use less energy, and vice-versa (Henderson et al., 2022; Bakhtiarifard et al., 2024).
-Additional results for these models can be found in App.
-G.
+Fig.
+5 summarises emissions for training (a), one-step inference (b), and rollout inference (c).
+Comparing total training emissions, the carbon footprints of the FNO variants are between 2.4 and 7.4 times lower than those of the U-net variants.
+For one-step inference the pattern changes.
+Here, the FNO variants are on average more expensive than the U-nets.
+During rollout inference, however, the differences between models become smaller.
+This is likely because the fixed overhead of model loading constitutes a larger fraction of the total cost.
+These discrepancies highlight the importance of considering both training and inference costs rather than focusing on a single stage of the modelling pipeline.
+Although carbon footprint and runtime are naturally correlated, they are not strictly proportional across models (Fig.
+6).
 5 Discussions
+Figure 6: Training runtime versus carbon footprint. The plot shows that energy consumption does not scale one-to-one with runtime.
+Fig.
+6 emphasises the importance of explicit energy or carbon measurement.
+Runtime alone is an unreliable proxy, as different hardware operations have distinct power profiles and a model that trains faster does not necessarily train cheaper [5].
 Our study demonstrates that physics inductive biases impact the trade-off between efficacy and efficiency.
-In simple settings, such as the harmonic oscillator, adding periodic structure or a physics-informed loss enables extrapolation beyond the training range.
-With viscous Burgers’ equation, combining data with physics yields the model that dominates the Pareto front, achieving lower test errors at lower carbon cost compared to the purely data-driven baseline.
+In the simple PINN experiment (Sec.
+1), combining data with physics yielded the model dominating the Pareto front, achieving lower test errors at lower carbon cost compared to both the purely data-driven and purely physics-driven baselines.
 In the incompressible shear flow setting, we observe clear differences in how the physics priors affect both predictive performance and carbon footprint across training and inference.
-Models with strong physics inductive biases achieve the lowest training footprints, but this comes at the expense of higher inference costs compared to U-net variants.
+Incorporating weak physics-inductive biases in the form of periodic padding, improves performance but increases total training emission, while its effects on training carbon efficiency and one-step inference are model-dependent: increasing both in one case (UNet-CP) and decreasing both in the other (UNet-CN-CP).
+Models with stronger physics inductive biases achieve the lowest training footprints, but this comes at the expense of higher inference costs compared to U-net variants.
 UFNO achieves one of the most favourable overall profiles, combining strong predictive accuracy with a low training footprint.
 Its inference cost, however, remains nearly twice that of the best-performing U-net variant (UNet-CN-CP).
-The results underscore the importance of evaluating training and inference together.
+The results underscore the importance of evaluating training and inference costs together.
 A model that trains cheaply may prove inefficient when repeatedly deployed, while others entail higher upfront cost but amortise more effectively in deployment.
-Amortising development costs across the model lifecycle is one way to factor in these considerations (Cottier et al., 2024).
-FM stands out as an outlier: while competitive in predictive performance, its reliance on numerical integration makes inference substantially more expensive.
-This is in spite of it operating on downsampled data, as the original resolution was computationally prohibitive (Dao et al., 2023).
+Amortising development costs across the model lifecycle is one way to factor in these considerations [33].
 Several key points follow from these results.
 First, model selection cannot be based on efficacy alone, as models with similar predictive accuracy can differ substantially in carbon emissions.
-Second, once both training and inference costs are accounted for, identifying a single “best” model becomes less straightforward, as illustrated by the incompressible shear flow experiment.
+Second, once both training and inference costs are accounted for, identifying a single “best” model becomes less straightforward.
 A model with a small carbon footprint during training may be unsuitable in deployment, making carbon-aware evaluation across the full lifecycle essential.
@@
 Many scientific and engineering problems are governed by well-understood structures, and overlooking them is wasteful.
-At the same time, tailoring models too closely to the specific physics can limit applicability and can be challenging for complex systems.
-This motivates the development of general-purpose models such as FNOs and FM, which allow physics priors to be incorporated more easily while retaining flexibility across tasks.
-However, as demonstrated by FM, alignment with the underlying physics does not guarantee carbon efficiency.
+At the same time, tailoring models too closely to the specific physics can limit applicability and can be challenging for complex systems (e.g., PINNs).
+This motivates the development of general-purpose models such as FNOs, which allow physics priors to be incorporated more easily while retaining flexibility across tasks.
+However, alignment with the underlying physics does not guarantee carbon efficiency.
 Physics should not be incorporated at any cost, but only where it provides clear benefits in terms of predictive performance or carbon efficiency.
-While our discussion centres on physics-inductive biases, the same logic extends to other forms of prior knowledge: any regularising bias that constrains the learning task appropriately can potentially offer similar trade-offs between efficacy and efficiency.
+While our work centres on physics-inductive biases, the same logic extends to other forms of prior knowledge: any regularising bias that constrains the learning task appropriately can potentially offer similar trade-offs between efficacy and efficiency.
 Ultimately, our results argue for a shift in mindset.
-Progress in ML should not be equated with ever-larger datasets, models, and compute budgets, but rather evaluated in terms of the value gained relative to the cost incurred.
+Progress in ML should not be equated with ever-larger datasets, models, and compute budgets, but rather evaluated in terms of the value gained relative to the cost incurred [34].
 Learning everything from data while ignoring domain knowledge reflects an abundance mindset, in which resources are treated as unlimited.
 A more sustainable alternative is to exploit what we already know by embedding this knowledge into models.
-Trading carbon for physics offers a principled path towards architectures that leverage prior knowledge, reduce reliance on brute-force compute, and achieve more favourable efficacy-efficiency trade-offs.
+This work offers a step towards understanding how prior knowledge shapes the efficacy-efficiency trade-offs, and argues for evaluating these explicitly.
 Limitations.
-This study has some limitations.
-First, all models were trained without extensive hyperparameter tuning.
-This choice emphasises architectural comparison over peak performance, trading statistical robustness for a broader perspective on performance–carbon trade-offs.
-In practice, however, hyperparameter search can dominate computational cost, and complex models typically require many more training runs during development than simpler baselines.
-Second, we restricted our study to spatio-temporal forecasting.
+We did not perform full hyperparameter sweeps for the models.
+As detailed in App.
+0.C, all configurations follow settings from prior work tuned on several PDE datasets, including the shear-flow dataset used here.
+This may influence the observed trade-offs.
+Furthermore, we restricted our study to spatio-temporal forecasting.
 While this limits the scope of the contribution, the central message about taking efficiency considerations into account by including appropriate inductive biases is applicable across ML tasks.
-Finally, our measurements do not account for hardware specific optimisations, where certain operations are more efficiently executed on specific accelerators.
-This, for instance, could yield an advantage to deep learning models compared to PDE solvers.
+Finally, the evaluation metrics focus on predictive accuracy rather than physical fidelity.
+Physics-based metrics would more directly assess whether the models capture underlying structures or conservation properties.
 6 Conclusions
 The primary message in this work, backed by empirical evidence, is to increase resource-awareness when designing ML experiments.
-While this can be accomplished in several ways (Bartoldson et al., 2023), we focus on the role of physics inductive biases and illustrate the different decisions practitioners have to consider.
-Through multiple experiments from simple oscillators to complex shear flow, we showed how physics-inductive biases affect the trade-off between predictive performance and carbon footprint.
+While this can be accomplished in several ways [35], we focus on the role of physics inductive biases and illustrate the different decisions practitioners have to consider.
+Using spatio- temporal forecasting of incompressible shear flow, we show how physics-inductive biases affect the trade-off between predictive performance and carbon footprint.
 This work advances a perspective grounded in resource-awareness and calls for a shift in mindset.
-As a community, we should discourage the pursuit of marginal improvements in performance that come from disproportionality higher resource costs.
+As a community, we should discourage the pursuit of marginal improvements in performance that come from disproportionately higher resource costs.
 With our work we have made the case to encourage resource-awareness by utilising existing domain knowledge and/or other inductive biases.
-Acknowledgements
+{credits}
+6.0.1 Acknowledgements
 SW and RS acknowledge funding received from Independent Research Fund Denmark (DFF) under grant agreement number 4307-00143B.
@@
 The authors thank members of SAINTS Lab for useful discussions throughout.
+[…]
+The authors have no competing interests to declare that are relevant to the content of this article.
 Appendix
-Appendix A Flow Matching
-We condition on two consecutive states $(u_{0},u_{1}):=(u_{k},u_{k+1})$ such that the path interpolates between current state $u_{0}$ and the target state $u_{1}$.
-Following Lim et al.
-(2025), we adopt their interpolation path:
-$p_{\tau}(u\mid u_{0},u_{1})=\mathcal{N}\!\left(u\;\middle|\;(1-\tau)u_{0}+\tau u_{1},\;\big(\sigma_{\min}^{2}+\sigma^{2}\tau(1-\tau)\big)I\right),$ (2)
-with $\sigma_{\min}=10^{-3}$ and $\sigma=10^{-1}$, as recommended for Navier–Stokes data.
-The mean interpolates linearly between $u_{0}$ and $u_{1}$, while the variance follows a Brownian-bridge profile (maximal halfway at $\tau=0.5$, minimal at endpoints at $\tau=0$ and $\tau=1$).
-The corresponding vector field is:
-$v_{\tau}(u\mid u_{0},u_{1})=u_{1}-u_{0}+\frac{\sigma^{2}}{2}\,\frac{1-2\tau}{\sigma_{\min}^{2}+\sigma^{2}\tau(1-\tau)}\Big(u-\big((1-\tau)u_{0}+\tau u_{1}\big)\Big).$ (3)
-The parametric vector field $v_{\tau}^{\theta}$ is trained to approximate $v_{\tau}$ by minimizing:
-$\mathcal{L}(\theta)=\mathbb{E}_{\tau,\;p_{\tau}(u\mid u_{0},u_{1}),\;q(u_{0},u_{1})}\Big\|v_{\tau}^{\theta}(u)-v_{\tau}(u\mid u_{0},u_{1})\Big\|^{2}.$ (4)
-Here $\tau\sim\mathcal{U}[0,1]$, $u\sim p_{\tau}(u\mid u_{0},u_{1})$, and $(u_{0},u_{1})\sim q(u_{0},u_{1})$, where $q$ denotes the empirical distribution of consecutive state pairs in the dataset.
-In addition to the conditioning pair $(u_{0},u_{1})$, we provide temporal context by sampling one additional state $u_{c}\in\{u_{k-3},u_{k-2},u_{k-1}\}$, giving a maximum horizon $h=4$ consistent with FNOs and U-nets.
-All inputs are downsampled by a factor of 16 via average pooling, and predictions are upsampled to the original resolution with bilinear interpolation.
-During inference, given a state $u_{k}$, the learned flow is simulated by numerically integrating $v_{\tau}^{\theta}$ from $\tau=0$ to $\tau=1$.
-We use Euler’s method with $10$ integration steps, which produces intermediate states approximating the continuous trajectory.
-The final iterate is then upsampled to the original resolution, yielding the forecast $\hat{u}_{k+1}$.
-Appendix B Toy Experiment on Harmonic Oscillator
-We design a toy experiment to study the effect of physics inductive biases for extrapolation in dynamical systems.
-The task is to model the trajectory of a harmonic oscillator, comparing standard feedforward networks with variants that incorporate different inductive biases.
-Specifically, we evaluate four networks with identical architecture but different activation functions, ReLU, Tanh, Sine, and Snake (Ziyin et al., 2020), where:
-$\text{Sine}(x)=\sin(x),\qquad\text{Snake}(x)=x+\sin^{2}(x).$ (5)
-The Sine and Snake activations introduce oscillatory biases aligned with the dynamics of the oscillator.
-In addition, we include a PINN that augments the data loss with a PDE residual while using Snake activations.
-The governing dynamics are given by the damped harmonic oscillator:
-$m\frac{d^{2}u}{dt^{2}}+\mu\frac{du}{dt}+ku=0\penalty 10000\ ,$ (6)
-$u(0)=1,\penalty 10000\ \penalty 10000\ \mathrm{and}\penalty 10000\ \penalty 10000\ \frac{du}{dt}=0.$ (7)
-The damping ratio is defined as $\delta=\mu/(2m)$.
-We adopt non-dimensionalized units with $m=1$ and $\omega_{0}=\sqrt{k/m}=1$.
-We consider both the undamped case ($\delta=0$) and a damped case ($\delta=0.02$).
-Training data are sampled from $t\in[-5\pi,5\pi]$ with Gaussian noise ($\mu=0$$\sigma=0.3$).
-The purely supervised networks use $10,000$ supervised samples, while the PINN uses $100$ points ($1\%$) for the physics loss.
-Test data are noise-free and sampled over $t\in[5\pi,25\pi]$ to assess extrapolation.
-The points for evaluating the physics-based loss are sampled across the entire domain.
-The purely supervised networks minimize the MSE:
-$\mathcal{L}_{\text{NN}}=\mathcal{L}_{\text{data}}=\frac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2},$ (8)
-where $u$ denotes the ground truth and $\hat{u}$ the model predictions.
-The PINN combines the same data term with a differential equation (DE) residual:
-$\mathcal{L}_{\text{PINN}}=\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{pde}}=\frac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2}+\frac{1}{N_{\text{de}}}\sum_{j=1}^{N_{\text{de}}}\left(\frac{d^{2}\hat{u}_{j}}{dt^{2}}+\mu\frac{d\hat{u}_{j}}{dt}+k\hat{u}_{j}\right)^{2}.$ (9)
-Further architectural and optimisation details are given in Sec. 4.1.
-Appendix C Loss Functions Used in the viscous Burgers’ Equation Experiment
-The loss function for the purely data-driven NN, denoted $\mathcal{L}_{\text{NN}}$, is the MSE loss given by:
-$\mathcal{L}_{\text{NN}}$ $=\mathcal{L}_{\text{data}}=\dfrac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2}.$ (10)
-Here $u$ and $\hat{u}$ denote the ground truth and the predictions.
-The U-PINN optimizes solely using the governing equations including those that describe the boundary and initial conditions to ensure a unique solution.
-Its loss function, $\mathcal{L}_{\text{U-PINN}}$, is composed of the residuals of the PDE and the deviations from the initial and boundary conditions, expressed as:
+Appendix 0.A Simple Experiment with PINNs
+Figure 7: Target and training data for the viscous Burgers’ equation. a: Target data. b: Supervised samples. c: PDE residual samples. d: Initial and boundary condition points.
+0.A.1 Experimental Setup
+In this simple experiment, we explore PINNs [36] which embed governing PDEs directly into the training objective, thereby regularising the optimisation landscape and guiding solutions towards physical consistency.
+By minimizing PDE residuals, PINNs can be trained with little or no supervised data.
+We study the solution of the viscous Burgers’ equation with Dirichlet boundary conditions on a 2D spatio-temporal grid:
+$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}=0,$ (2)
+$u(x,0)=-\sin(\pi x),$ (3)
+$u(1,t)=u(-1,t)=0,$ (4)
+for $x\in[-1,1]$, $t\in[0,1]$, and with fixed viscosity ($\nu=0.025$). The domain is discretized with $256$ spatial points and $100$ time steps (Fig. 7).
+We compare three models that share the same underlying architecture (four hidden layers, width 50, tanh activations) but differ in their use of physics through the loss function.
+The purely data-driven NN minimizes the mean-squared-error (MSE) between predictions $\hat{u}$ and ground truth $u$:
+$\mathcal{L}_{\text{NN}}=\mathcal{L}_{\text{data}}=\frac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2}.$ (5)
+The unsupervised U-PINN is trained solely on physics constraints, minimizing PDE residuals alongside deviations from initial and boundary conditions:
 $\mathcal{L}_{\text{U-PINN}}$ $=\mathcal{L}_{\text{pde}}+\mathcal{L}_{\text{ic}}+\mathcal{L}_{\text{bc}}$ (6)
-$=\frac{1}{N_{\text{pde}}}\sum_{i=1}^{N_{\text{pde}}}\left(\frac{\partial\hat{u}_{i}}{\partial t}+\hat{u}_{i}\frac{\partial\hat{u}_{i}}{\partial x}-\nu\frac{\partial^{2}\hat{u}_{i}}{\partial x^{2}}\right)^{2}$ (12)
+$=\frac{1}{N_{\text{pde}}}\sum_{i=1}^{N_{\text{pde}}}\left(\frac{\partial\hat{u}_{i}}{\partial t}+\hat{u}_{i}\frac{\partial\hat{u}_{i}}{\partial x}-\nu\frac{\partial^{2}\hat{u}_{i}}{\partial x^{2}}\right)^{2}$
 $+\frac{1}{N_{\text{ic}}}\sum_{j=1}^{N_{\text{ic}}}\left(\hat{u}(x_{j},0)+\sin(\pi x_{j})\right)^{2}$
 $+\frac{1}{N_{\text{bc}}}\sum_{k=1}^{N_{\text{bc}}}\left(\hat{u}(-1,t_{k})^{2}+\hat{u}(1,t_{k})^{2}\right).$
-The S-PINN combines both data and physics losses. Its total loss function, $\mathcal{L}_{\text{S-PINN}}$ is given by:
-$\mathcal{L}_{\text{S-PINN}}=\mathcal{L}_{\text{NN}}+\mathcal{L}_{\text{U-PINN}}=\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{pde}}+\mathcal{L}_{\text{ic}}+\mathcal{L}_{\text{bc}}.$ (13)
-Appendix D Incompressible Shear Flow Dataset
-This experiment is based on data from The Well (Ohana et al., 2024), consisting of 2D incompressible Navier–Stokes dynamics with an additional passive tracer:
+The semi-supervised S-PINN combines both objectives, with data and physics terms weighted equally:
+$\mathcal{L}_{\text{S-PINN}}=\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{pde}}+\mathcal{L}_{\text{ic}}+\mathcal{L}_{\text{bc}}.$ (7)
+All models were trained using Adam optimizer followed by L-BFGS optimizer [37] with learning rate $10^{-3}$.
+Dataset sizes range from $10^{3}$ to $5\times 10^{3}$ training points for PINNs and up to $8\times 10^{3}$ for the NN.
+Data and code are adapted from [38].
+0.A.2 Results
+Both PINNs achieve consistently lower test losses at lower carbon footprints compared to the supervised NN (Fig.
+1-b).
+While the NN has the lowest carbon footprint for a fixed number of epochs and data points, achieving a given accuracy level is up to $10$ times more carbon-efficient with the PINNs.
+The S-PINN, which combines data and physics, dominates the Pareto front, suggesting that the right balance between data and physics yields the most favourable trade-off between predictive performance and carbon cost.
+Appendix 0.B Incompressible Shear Flow Dataset
+This experiment is based on data from The Well [26], consisting of 2D incompressible Navier–Stokes dynamics with an additional passive tracer:
 $\partial_{t}\mathbf{u}+\nabla p-\nu\nabla^{2}\mathbf{u}$ $=-\mathbf{u}\cdot\nabla\mathbf{u},$ (8)
@@
 Thus, the initial conditions are fully characterized by three parameters: $n_{\text{shear}}\in\{2,4\}$, $n_{\text{blobs}}\in\{2,3,4,5\}$, and $w\in\{0.25,0.5,1.0,2.0,4.0\}$.
-An example of the data is illustrated in Fig.
-9.
-As part of preprocessing, we standardize each field using $\hat{x}=(x-\mu)/\sigma$ (see Table 2).
+As part of preprocessing, we standardize each field using $\hat{x}=(x-\mu)/\sigma$.
 This normalization balances the contribution of different fields by preventing those with high variance from dominating the loss.
-Table 2: Field statistics. Mean and standard deviation of raw fields, averaged over all time steps.
-Field Mean Std
-s $2.19\times 10^{-3}$ $0.33$
-p $1.18\times 10^{-9}$ $0.074$
-$\mathbf{v_{x}}$ $2.19\times 10^{-3}$ $0.40$
-$\mathbf{v_{y}}$ $1.40\times 10^{-7}$ $0.097$
-Figure 9: Example of incompressible shear flow data.
-Temporal evolution of tracer (s), pressure (p), horizontal velocity ($v_{x}$) and vertical velocity ($v_{y}$) fields at four time steps: $t\in\{0,T/3,2T/3,T\}$.
-The snapshots illustrate the coupled dynamics of the fields.
-Appendix E Flow Matching in Low-Resolution Data Space
-Most prior work applies FM in latent space, with a few exceptions (Disch et al., 2025; Baldan et al., 2025) demonstrating its feasibility directly in data space.
-We adopt this latter strategy to avoid the additional energy and carbon costs of an encoder–decoder pipeline.
-To make FM computationally tractable, we downsample inputs by a factor of 16, perform training and inference at low resolution, and subsequently upsample to the original resolution.
-This design inevitably encode a trade-off between efficacy and efficiency: while operating in low-resolution data space suppresses fine-scale dynamics and may introduce artifacts, it delivers substantial efficiency gains, reducing training and inference time by orders of magnitude.
-Appendix F Model Configurations
+Appendix 0.C Model Configurations
 Model configurations are adapted from prior work.
 U-nets, FNO, and TFNO follow Ohana et al.
-(2024), UFNO follows Rahman et al.
-(2023), and FM is adapted from Lim et al.
-(2025), with some adjustments to match our setting (including operating directly in data space).
-In line with The Well (Ohana et al., 2024), U-nets and FNO variants are scaled to $\sim$17–20M parameters, while FM is smaller ($\sim$8.5M).
-Table 3 provides an overview of parameter counts, with detailed configurations summarized in Table 4 (U-nets and FNOs) and Table 5 (FM).
-Table 3: Model parameters.
-Model #Params
-UNet 17.517.5M
-UNet-CN 18.618.6M
-UNet-CP 17.517.5M
-UNet-CN-CP 18.618.6M
-FM 8.68.6M
-FNO 19.019.0M
-TFNO 19.319.3M
-UFNO 18.818.8M
-Table 4: U-nets and FNOs model configurations.
-Model Filter size/modes Init dim Blocks/stage Up/down Bottleneck
-UNet, UNet-CP 33 $48$ 11 44 11
-UNet-CN, UNet-CN-CP 77 $42$ 22 44 11
-FNO, TFNO $16$ $128$ 44
-UFNO $8,12,16$ $128$ 11 33 11
-Table 5: FM model configuration. The FM model operates in low-resolution data space and is built from transformer stacks with learned positional, timestamp, and distance embeddings.
-Model Resolution Init dim Depth Bottleneck Heads
-FM $16\times 32$ $256$ 33 44 88
-Appendix G Additional Results
-This appendix presents supplementary plots that support the findings in the main text. Table 6 reports both the training and inference emissions alongside run times.
-Training Inference (one-step)
-Model CO2eq (kg) Time (h) CO2eq (g) Time (min)
-UNet $2.72$ $12.7$ $3.68\pm 0.03$ $1.0$
-UNet-CN $4.71$ $15.9$ $7.77\pm 0.05$ $1.9$
-UNet-CP $3.81$ $14.2$ $4.06\pm 0.05$ $1.0$
-UNet-CN-CP $5.89$ $20.9$ $4.87\pm 0.02$ $1.2$
-FM $3.31$ $14.0$ $23.54\pm 0.13$ $4.6$
-FNO $0.80$ $4.0$ $7.87\pm 0.94$ $2.0$
-TFNO $1.09$ $5.3$ $7.21\pm 0.12$ $1.8$
-UFNO $1.12$ $3.6$ $9.00\pm 0.88$ $2.4$
-Table 6: Carbon footprint during training and inference.
-Reported CO2eq emissions and runtime for all models, measured with CarbonTracker.
-Inference results correspond to one-step forecasts over the full test set (4,704 samples) and includes the initial model loading cost.
-Each inference experiment was repeated three times; CO2eq is reported as mean and standard deviation, and runtimes as mean values.
-Fig.10-a reports model performance using the variance-scaled root mean squared error (VRMSE), which normalises prediction error by the variance of the ground truth to account for scale differences across fields:
-$\text{VRMSE}(u,\hat{u})=\left(\frac{\langle|u-\hat{u}|^{2}\rangle}{\langle|u-\langle u\rangle|^{2}\rangle+\epsilon}\right)^{1/2},$ (19)
-where $\langle\cdot\rangle$ denotes spatial averaging, $u$ the ground truth, $\hat{u}$ the prediction, and $\epsilon=10^{-8}$ a small constant for stability.
-The VRMSE measures errors relative to the natural variability of the data, with $\text{VRMSE}(y,\langle y\rangle)\approx 1$ implying that values greater than one indicate prediction errors larger than typical fluctuations.
-The results follow the same overall trends as in Fig.8-a, where performance was measured with the Pearson correlation coefficient.
-Figure 10-b reports the inference carbon footprint for $24$ autoregressive rollouts of $50$ steps each.
-Compared to one-step inference over the full test set ($4,704$ samples), the rollout costs tend to average out across models, since the fixed overhead of model loading constitutes a larger share of the total cost.
-Flow matching remains an outlier, standing out as the most expensive model during inference.
-Figure 10: Predictive performance and rollout carbon footprint.
-a: VRMSE vs.
-training carbon footprint.
-Each point shows every second step of a 20-step rollout, averaged across the four fields (lighter shades indicate earlier steps).
-b: Carbon footprint of inference during rollouts.
-Reported CO2eq emissions for 24 autoregressive rollouts of 50 steps.
-Inference costs include the initial model loading cost.
-Models are ordered from no bias (left) to strong bias (right).
+[26], while UFNO follows Rahman et al.
+[30].
+All models are scaled to 17–20M parameters.
+Table 3 summarises configurations and parameter counts.
+Table 3: Model configurations.
+Model Filter size/modes Init dim Blocks/stage Up/down Bottleneck Params ($\times 10^{6}$)
+UNet, UNet-CP 33 $48$ 11 44 11 $17.5$
+UNet-CN, UNet-CN-CP 77 $42$ 22 44 11 $18.6$
+FNO, TFNO $16$ $128$ 44 $19.0$, $19.3$
+UFNO $8,12,16$ $128$ 11 33 11 $18.8$
+Appendix 0.D Rollout Results
 Figs.
-11, 12, and 13 provide rollout prediction examples for the three physical fields not shown in the main text (pressure fields and the two velocity components).
-The observations made for the tracer field in the main text extend to these additional fields.
-Figure 11: Example of rollout predictions for pressure field.
-Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout.
-Models labels are ordered from no bias (top) to strong bias (bottom).
+8-11 provide rollout prediction examples for each of the four fields.
+The baseline U-net (UNet and UNet-CP) deteriorate rapidly, accumulating distortions within just a few steps.
+The modernised variants (UNet-CN, UNet-CN-CP) show marked improvements: UNet-CN captures fine-scale detail but gradually disrupts large-scale dynamics, while UNet-CN-CP better preserves structure at both scales at the cost of high-frequency artifacts.
+FNO and TFNO maintain dominant large-scale dynamics but exhibit a systematic drift and fail to capture fine-scale variability.
+UFNO recovers both large- and small-scale structures more effectively, though these gradually smooth out over time.
 Figure 8: Example of rollout predictions for horizontal velocity field.
@@
 Models labels are ordered from no bias (top) to strong bias (bottom).
+Figure 10: Example of rollout predictions for pressure field.
+Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout.
+Models labels are ordered from no bias (top) to strong bias (bottom).
+Figure 11: Example rollout predictions for the tracer field.
+Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout.
+Models labels are ordered from no bias (top) to strong bias (bottom).